### v1

# CAPE: A CLIP-Aware Pointing Ensemble of Complementary Heatmap Cues for Embodied Reference Understanding

Abstract
We address the problem of Embodied Reference Understanding, which involves predicting the object that a person in the scene is referring to through both pointing gesture and language. Accurately identifying the referent requires multimodal understanding: integrating textual instructions, visual pointing, and scene context. However, existing methods often struggle to effectively leverage visual clues for disambiguation. We also observe that, while the referent is often aligned with the head-to-fingertip line, it occasionally aligns more closely with the wrist-to-fingertip line. Therefore, relying on a single line assumption can be overly simplistic and may lead to suboptimal performance. To address this, we propose a dual-model framework, where one model learns from the head-to-fingertip direction and the other from the wrist-to-fingertip direction. We further introduce a Gaussian ray heatmap representation of these lines and use them as input to provide a strong supervisory signal that encourages the model to better attend to pointing cues. To combine the strengths of both models, we present the CLIP-Aware Pointing Ensemble module, which performs a hybrid ensemble based on CLIP features. Additionally, we propose an object center prediction head as an auxiliary task to further enhance referent localization. We validate our approach through extensive experiments and analysis on the benchmark YouRefIt dataset, achieving an improvement of approximately 4 mAP at the 0.25 IoU threshold.
1 Introduction
Embodied Reference Understanding (ERU) [11] is the task of identifying a specific object in a visual scene based on language instructions and pointing cues in the image. It plays a key role in real-world applications like human-robot interaction, assistive robotics, and augmented reality, where systems must understand which object a person refers to.
While grounding models [37] have made significant progress in detecting objects mentioned in natural language, they often fall short in ERU task, particularly in ambiguous scenes. When multiple instances of the same object type are present, these models tend to detect all matching candidates without the ability to disambiguate and identify the specific target intended by the user. Moreover, when the textual instruction itself is vague or ambiguous, grounding models struggle even further, often failing to identify the correct target object or producing no confident prediction at all due to the inherent ambiguity. This limitation highlights the need for additional disambiguation cues, embodied gesture signals, that can help resolve referential ambiguity and enable accurate identification of the intended object. Therefore, both embodied gesture signals and language reference are crucial to identify the referent. In ERU, two main challenges arise. (1) identifying candidate objects from the text, as in standard grounding tasks, and (2) interpreting the pointing cue from visual input to guide the final prediction. The latter requires understanding human pose, inferring pointing direction, and handling visual complexities like perspective, occlusion, and depth.
In [11], first time ERU task is defined and a multimodal YouRefIt dataset is proposed. They also proposed a method that combines textual and visual inputs along with PAF [7] and predicted saliency maps [33] to help localize the human and narrow down the potential referent. While these inputs lead to improved performance, the model still exhibits suboptimal accuracy in many cases, particularly due to its inability to properly follow the pointing instructions. Building on this, the Touch-Line Transformer [35] introduces a novel approach by predicting a Virtual Touch-Line and applying losses that encourage the model to focus on the head-to-fingertip direction, which they find often aligns closely with the intended referent [24, 44]. Although this method achieves significant gains, it still faces challenges. The model occasionally overlooks the pointing line, as it is not explicitly provided as input. Instead, it predicts head, fingertips, and arm position to encourage attention in that direction. More importantly, the head-to-fingertip direction does not always align with the target. In some scenarios, the accurate pointing direction follows a line from the hand alone (e.g., wrist-to-fingertip), rather than from the head (Fig. 1). These observations suggest that while head-to-fingertip direction offers valuable guidance, it is not sufficient on its own to fully resolve the complexity of the task.
To tackle these challenges, we propose a dual-model framework that leverages complementary pointing cues. Both models take as input a referring expression and an image with a pointing gesture, along with an additional heatmap that encodes the pointing direction. One model utilizes a head-to-fingertip heatmap, while the other uses a wrist-to-fingertip heatmap, capturing different aspects of the pointing behavior. Their predictions are integrated via a CLIP-Aware Pointing Ensemble (CAPE) module, which effectively combines the strengths of both models. Additionally, we introduce an auxiliary objective to predict the object center, providing an extra supervisory signal that further enhances grounding accuracy. Our contributions are: (1) We propose using heatmap as extra modality to guide our end-to-end reference object detection model to focus on pointing information more. (2) We propose two parallel complementary models to focus on head-to-fingertip and wrist-to-fingertip pointing lines, to individually alleviate inherent challenges of the task. (3) We present CAPE module to effectively ensemble two complementary models. (4) We introduce object center prediction head to enhance prediction accuracy by focusing on finding the object without considering the precise bounding box.
2 Related Work
Referring Expression Comprehension (REC)
In REC [23, 31, 42, 46, 47, 72], the goal is to identify a specific region in an image based on a given referring expression. In contrast to traditional object detection, REC interprets free-form text and can locate objects from any category, including previously unseen ones. Similarly, OV-DETR [73] integrates image and text embeddings from a CLIP model as queries within the DETR [9] to generate category-specific bounding boxes. ViLD [19] distills knowledge from a CLIP teacher model into an R-CNN-like detector [18], enabling region embeddings to capture semantic information from language. GLIP [17] formulates object detection as a grounding task, using additional grounding datasets to align visual regions with textual phrases. DetCLIP [71] enriches its knowledge base using generated pseudo labels. While YOLO-World [12] extends traditional YOLO object detection [50] to an open-world setting, GroundingDINO [37] aligns visual and textual features to detect arbitrary objects described by free-form text. Recently, LMMs [5, 1, 3] get attention for their superior performance in visual grounding [10].
Nonverbal Communication for Referent
Some studies [55, 56, 57, 67, 6, 66, 25, 45, 61, 29, 53, 64, 63, 20, 60, 68, 58, 62, 59, 21, 51] use gaze target detection as a nonverbal cue, however, gaze alone is unreliable due to distractions and the absence of a clear pointing moment. It is typically localized frame by frame and often used to support conversation [4, 8]. In [43], a multimodal exophora resolution method is proposed to disambiguate demonstrative expressions like “that one” by integrating object categories, pointing gestures, and prior environmental knowledge. [14] introduced an interactive robot dialogue system that uses multimodal interaction and pointing line estimation to accurately identify referent with an iterative correction process. Their follow-up work [15] extended this by recognizing unseen objects using a Region Proposal Network and VL-T5 multimodal network [13], moving beyond general object classes from benchmarks like MSCOCO [36] to attribute-based object recognition. On the other hand, several datasets have been collected to incorporate nonverbal cues, but most remain in simulated domains due to accessibility and controllability [26, 27, 2, 28].
Embodied Reference Understanding (ERU)
ERU is a recent advancement of REC that considers the subject’s position while pointing at a reference object. Chen et al. [11] introduced the task and benchmarks using both verbal (text) and nonverbal (pointing gesture) cues. Their model leverages predicted saliency maps and PAF [7] as gestural features to better perceive pointing direction and detect referents. To address ambiguities caused by camera perspective, [52] mapped scenes to 3D coordinates using depth estimation and subject position for spatial attention, but this only yielded modest improvements. Li et al. [35] further improved ERU using an MDETR-based [30] model, introducing a “virtual-touch-line” from eye to fingertip and predicting its vector alongside bounding boxes. In this work, we also use a transformer-based multimodal object detector similar to MDETR [30]. However, instead of predicting a single line vector, we extend the visual-touch-line concept to a pointing heatmap representing the focus of attention. To capture different pointing variations, we use two different heatmaps processed in parallel networks. Some recent works [39, 40] have extended ERU task to 3D embodied settings. ScanERU [39] introduced the first 3D-ERU benchmark by inserting human avatars into existing 3D datasets, but relied on manual placement. In contrast, Ges3ViG [40] improves realism and scalability by automating avatar insertion, generating gesture-aware instructions, and incorporating human localization into the grounding pipeline.
3 Methodology
Problem Definition In ERU, given an RGB image and a text input , we predict a bounding box corresponding to an object referenced by the text and indicated by a pointing gesture in the image. We propose using an additional pointing heatmap to supervise the model better focus on the visual pointing reference.
3.1 Model Architecture
Fig. 2 illustrates our architecture, which consists of a large-scale text encoder, image encoder, and heatmap encoder. The outputs of these encoders are concatenated and fed into a transformer [65] encoder block. Subsequently, a transformer decoder processes the encoder output to generate predictions through task-specific heads. In this work, we introduce two networks that share the same architecture but are trained independently. The first network () receives a heatmap generated from the wrist-to-fingertip line, while the second () uses a heatmap derived from the head-to-fingertip line. Finally, a CLIP-Aware Pointing Ensemble (CAPE) module combines the outputs of these two complementary models to enhance overall performance.
Encoders
We use a pretrained ROBERTA [38] which is a robustly optimized version of BERT [16] as the text encoder to obtain textual embeddings from the input text. To obtain image embeddings, , we utilize a pretrained ResNet-101 [22]. Since the pointing heatmap () is a sparse representation, we choose a lightweight encoder, ResNet-18 [22], to embed it: This choice not only improves the computational efficiency of the network but also helps prevent redundant information in the final concatenated feature representation, which could otherwise harm the model’s learning capacity.
Transformer encoder
After obtaining the embeddings from three encoders, we first apply a convolution to each embedding individually to project their channel dimensions to . Next, we flatten the spatial dimensions of the embedding into a single dimension, converting them into sequences of tokens. For example, for the heatmap embedding: . We then concatenate the three embeddings at the sequence level to form the final representation, which is fed into the transformer encoder. It processes this concatenated input to learn multimodal representations.
Transformer decoder
We feed the multimodal representation output from the transformer encoder into the transformer decoder. Additionally, we provide a set of learnable object queries and gestural keypoint queries. The transformer decoder is responsible for generating object output embeddings and gestural output embeddings.
Prediction head
The object and gestural output embeddings produced by the transformer decoder serve as inputs to our prediction heads. These heads are responsible for predicting candidate bounding boxes for the referent, gestural keypoints (pointing direction prediction head in Fig. 2), which predicts eye, fingertip, wrist coordinates as well as arm class, and center points of the referent. Finally, we select one bounding box prediction, one center point and a pair of gestural keypoints with the highest confidence scores as the final prediction. Please note that we use multi-layer perceptrons (MLPs) as the prediction heads.
3.2 Pointing Heatmap Modality Learning
Incorporating a heatmap that represents the pointing direction provides a valuable spatial priors for the network, especially considering the fact that the existing models in this field tend to give less attention to the visual pointing. This explicit encoding pointing direction helps localize the target region by guiding the model’s attention toward direction and areas that are more likely to be referenced. This is particularly beneficial in real-world, cluttered scenes where gesture interpretation is inherently ambiguous.
How should we create heatmap?
The VTL [35] demonstrates superior performance in most cases, as it effectively captures the intended pointing direction by leveraging both the fingertip position and the head orientation. Human pointing gestures are typically guided by visual attention: we often align our finger with the target object based on our eye-level perspective [35]. Consequently, the head–to-fingertip line serves as a strong proxy for the true pointing intent, integrating both gaze and gesture cues. However, the VTL is not universally reliable. In certain scenarios, the performer may be looking elsewhere (for instance, toward a robot or camera) rather than at the intended object. Additionally, in cases involving extreme viewing angles or occluded head poses, the head orientation may become ambiguous or misleading. In such situations, the local orientation of the hand, as approximated by the wrist–to-fingertip line, can offer complementary directional cues that help disambiguate the pointing gesture. We hypothesize that heatmaps derived from the head-to-fingertip and wrist-to-fingertip lines provide complementary information. Therefore, we generate heatmaps that incorporate both information. For the head-to-fingertip heatmap (), we use the eye coordinates as the starting point and draw a line toward the edge of the image, passing through the fingertip. Similarly, for the wrist-to-fingertip heatmap (), we start from the wrist and draw a line to the edge of the image through the fingertip point. To generate these heatmaps, we apply a Gaussian Ray Heatmap (See Appendix B).
How should we integrate these heatmaps?
The next important question is how to integrate these heatmaps effectively. We explore two different approaches. In the first approach, we merge the two heatmaps () and feed the combined input into a single network to examine whether the model can implicitly learn to utilize both sources of information and determine the most appropriate direction based on the human pose. In the second approach, we train two identical networks independently and ensemble them: one using the head-to-finger heatmap () and the other using the wrist-to-fingertip heatmap (). We empirically find that the second approach yields better performance. As expected, it allows us to train two complementary networks. Notably, the network trained with the head-to-fingertip heatmap () demonstrates significant improvements. In contrast, the first approach results in only subtle enhancement, and the model fails to effectively leverage both sources of information to determine the pointing direction (see Tab. 3 Setup G). Its behavior closely resembles the head-to-fingertip-only scenario. However, due to the presence of additional wrist-to-fingertip line in the merged heatmap, the model experiences contradictory information, ultimately leading to reduced accuracy than head-to-fingertip heatmap-only setup.
Where should we inject these heatmaps?
We empirically find that the most effective way to leverage heatmaps is to use a separate encoder to extract heatmap embeddings, which are then concatenated with the output of the image encoder. It enables us to preserve the pretrained image encoder and allows the dedicated heatmap encoder to extract rich, modality-specific information.
3.3 CLIP-Aware Pointing Ensemble (CAPE)
We introduce an inference time ensemble module to increase the performance by benefiting from two complementary models, and . CAPE is an adaptive scoring method that combines model confidence and CLIP-based similarity in a size-aware manner effectively. (1) For each prediction (top-2 predictions from both models), we compute a normalized CLIP score [48]. We sum pointing models’ confidece scores with the CLIP scores to obtain the fused score (CLIP Fusion in Tab. 5). (2) We calculate CLIP scores for the top-1 predictions of both models. If the confidence scores of the second-highest predictions exceed a threshold, we also compute their CLIP scores. Finally, we select the prediction with the highest CLIP score from among these candidate boxes (CLIP-Only Top-2 + Threshold=0.95 in Tab. 5). CAPE applies strategy (1) when the referent is a small object (defined as occupying less than 0.48% of the image area, following [11]) to avoid relying solely on the CLIP prediction, as CLIP becomes less reliable for smaller objects. For all other objects, strategy (2) is used (see Sec. C.5 for details). This hybrid approach leverages the strengths of both signals while adapting to object scale. All thresholds and selection rules were tuned on the validation set to avoid test-time overfitting.
3.4 Gestural Signal Learning
Referent alignment loss
Following [35], we incorporate a referent alignment loss () to enforce consistency between the predicted referent and the VTL. The core idea is that a correct referent should be geometrically aligned with the pointing direction. To quantify this alignment, we compute the cosine similarity between the eye-to-fingertip vector and the eye-to-object vector, defined as:
where () and () are the eye and the fingertip coordinates. () are the center of the predicted bounding box. We compute the same similarity using GT bounding box to obtain . The referent alignment loss then penalizes deviations between the predicted and GT alignment:
Please note that for both and , we use GT eye and fingertip coordinates to ensure accurate directional representation. We apply this loss during the training of the model , which uses the head-to-fingertip heatmap as reference. For training the model, which utilizes the wrist-to-fingertip heatmap, we replace the eye coordinates with wrist coordinates. The rest remains unchanged, as we similarly aim to maximize the directional correlation.
Object center prediction
In our task, identifying the correct referent and accurately localizing it are two distinct but equally important challenges. To enhance the localization capability of the model, we decouple the prediction of the object center from the rest of the bounding box regression. Specifically, we introduce an additional prediction head implemented as a MLP at the end of the network. This head is dedicated to predicting the () coordinates of the referent object’s center, independently from the standard bounding box regression head. During training, we supervise this center prediction using the L1 loss between the predicted and GT center coordinates:
This explicit supervision encourages the model to focus on spatial alignment and leads to more accurate object localization, especially in scenes with dense or overlapping objects. Moreover, this center prediction serves as an auxiliary task, which improves the representational capacity of the model by encouraging the visual backbone to learn richer geometric and spatial features. We apply this auxiliary head and its corresponding loss to both models. Since this head is used only during training, it introduces no additional computational overhead during inference.
Gesture prediction
In addition to the bounding box and object center predictions, our model incorporates several auxiliary heads to predict task-specific features, including eye (wrist in ) and fingertip coordinates, and arm classification (See prediction head in Sec. 3.1). For each prediction, we apply a standard L1 loss between the predicted and GT data, which encourages spatial attention to task-relevant regions such as the eyes and fingertips, enhancing the model’s understanding of referential gestures. Moreover, we apply cross-entropy loss to classify whether the predicted gestural keypoints are correct.
Total loss
We define our objective as follows:
where is bounding box loss and it is the combination of L1 and GIoU losses as in DETR-like methods [9, 34, 41, 74, 75, 37]. denotes referent alignment loss, indicates object center prediction loss, and is gesture prediction loss. and are soft token loss and contrastive loss respectively to help visual and textural signals alignment as in [30]. We empirically determine the optimal coefficients on the validation set (See Appendix E).
| IoU Threshold for mAP | 0.25 | 0.50 | 0.75 | |||||||||
| Object Sizes | All | S | M | L | All | S | M | L | All | S | M | L |
| Grounding DINO [37] | 57.9 | 38.0 | 60.9 | 74.9 | 54.9 | 35.7 | 59.3 | 69.6 | 42.3 | 22.7 | 45.9 | 58.4 |
| FAOA [69] | 44.5 | 30.6 | 48.6 | 54.1 | 30.4 | 15.8 | 36.5 | 39.3 | 8.5 | 1.4 | 9.6 | 14.4 |
| ReSC [70] | 49.2 | 32.3 | 54.7 | 60.1 | 34.9 | 14.1 | 42.5 | 47.7 | 10.5 | 0.2 | 10.6 | 20.1 |
| YourRefit PAF [11] | 52.6 | 35.9 | 60.5 | 61.4 | 37.6 | 14.6 | 49.1 | 49.1 | 12.7 | 1.0 | 16.5 | 20.5 |
| YourRefit Full [11] | 54.7 | 38.5 | 64.1 | 61.6 | 40.5 | 16.3 | 54.4 | 51.1 | 14.0 | 1.2 | 17.2 | 23.3 |
| REP [52] | 58.8 | 44.7 | 68.9 | 63.2 | 45.7 | 25.4 | 57.7 | 54.3 | 18.8 | 3.8 | 22.2 | 29.9 |
| Touch-Line-EWL [35] | 69.5 | 56.6 | 71.7 | 80.0 | 60.7 | 44.4 | 66.2 | 71.2 | 35.5 | 11.8 | 38.9 | 55.0 |
| Touch-Line-VTL [35] | 71.1 | 55.9 | 75.5 | 81.7 | 63.5 | 47.0 | 70.2 | 73.1 | 39.0 | 13.4 | 45.2 | 57.8 |
| Ours (CAPE) | 75.0 | 63.2 | 80.2 | 81.8 | 65.4 | 49.5 | 74.3 | 72.7 | 35.7 | 13.4 | 40.1 | 53.5 |
4 Experimental Results
Dataset
For both training and testing, we use the image subset of the YouRefIt dataset [11], as the video portion is not publicly available. The dataset contains 2,950 training images and 1,245 test images. Annotations for bounding boxes, as well as eye, fingertip, elbow, and wrist coordinates, are provided by [11, 35].
Evaluation
For fair evaluation and comparison, we follow prior work [11] and report mean Average Precision (mAP) under three Intersection-over-Union (IoU) thresholds: 0.25, 0.50, and 0.75. A prediction is considered correct if its IoU with the GT exceeds the given threshold. Additionally, by following the literature, mAP scores are reported with respect to object size, categorized as Small (S), Medium (M), and Large (L). The object size categories are determined based on the ratio of the object area to the image area, using two thresholds [11]: and . An object is classified as small if its area ratio is below , medium if it falls between and , and large otherwise. In addition to this standard evaluation protocol, we introduce two metrics for assessing referential grounding performance: CLIP score and center coordinate distance (). To compute the CLIP score, we extract features from the text input and the cropped region corresponding to the predicted bounding box using CLIP’s text and image encoders. We then calculate the cosine similarity between these feature vectors to quantify the semantic alignment between the predicted object and the text. For , we compute the L1 distance between the center coordinates of the predicted bounding box and those of the GT bounding box, providing a direct measure of spatial alignment.
Implementation
We use the AMSGrad optimizer [49, 32] during training and train our models for 30 epochs. Both the transformer encoder and decoder consist of layers, each with attention heads and an MLP dimension of . We apply dropout [54] with in every layer of both the transformer encoder and decoder. The learning rate is set to for the text encoder and for the remaining components. All experiments are conducted on a single NVIDIA RTX A6000 GPU with a batch size of 4.
| CLIP Score | ||||||||
| Objet Sizes | All | S | M | L | All | S | M | L |
| Touch-Line-EWL | 0.7369 | 0.6936 | 0.7305 | 0.7845 | 0.3168 | 0.3006 | 0.2903 | 0.3564 |
| Touch-Line-VTL | 0.7368 | 0.6925 | 0.7321 | 0.7846 | 0.2843 | 0.2809 | 0.2276 | 0.3393 |
| Ours | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
4.1 Results
Tab. 1 compares the performance of our method with existing approaches on the YouRefIt benchmark. Our model achieves SOTA results across all objects at IoU thresholds of 0.25 and 0.50, despite being trained for only 30 epochs on a single GPU with a batch size of 4. According to our experiments, this limited training setup affects bounding box precision, contributing to the suboptimal performance at , alongside limitations related to the backbone (see Grounding DINO in Tab. 1 for backbone). In contrast, Touch-Line [35] was trained for 200 epochs with a batch size of 56. This clearly demonstrates the efficiency and effectiveness of our approach. Tab. 2 presents a comparison between our method and the previous SOTA, Touch-Line, using two new metrics in this field: CLIP score and . The higher CLIP scores indicate that our predicted bounding boxes have stronger semantic alignment with the input text, suggesting better object identification and improved bounding box precision. Similarly, our method achieves the best performance under the metric, which quantifies the distance between the centers of the predicted and GT bounding boxes. This confirms that our model produces more spatially accurate predictions. In Fig. 3, we present samples images with predicted and GT bounding boxes. We compare our model with Touch-Line [35].
| Setup | Method | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| A | Baseline | 71.2 | 60.1 | 32.8 | 0.7409 | 0.2662 |
| B | A + object center prediction | 70.8 | 61.6 | 34.6 | 0.7339 | 0.2707 |
| C | A + W2F heatmap | 68.9 | 59.7 | 32.4 | 0.7353 | 0.2984 |
| D | A + H2F heatmap | 71.9 | 62.8 | 33.8 | 0.7375 | 0.2694 |
| E | B + W2F heatmap () | 69.6 | 60.7 | 31.5 | 0.7344 | 0.2770 |
| F | B + H2F heatmap () | 72.9 | 62.3 | 35.1 | 0.7373 | 0.2490 |
| G | A + W2F heatmap + H2F heatmap | 70.2 | 60.2 | 33.8 | 0.7345 | 0.2744 |
| H | Full model - Ensemble of E and F | 75.0 | 65.4 | 35.7 | 0.7983 | 0.2476 |
4.2 Analysis
In this section, we provide a detailed analysis of the optimality of our design choices, including heatmap generation and the ensemble method. Specifically, we examine various heatmap injection strategies and ensemble approaches, among other factors. The baseline in Tab. 3 (Setup A) refers to a straightforward training of our model without incorporating any of the proposed contributions.
Object center prediction
The object center prediction head and its associated loss yield improvements over the baseline in terms of mAP at and . However, they have a detrimental effect on mAP at , as well as on the CLIP Score and . Notably, when combined with the heatmap (Setup E and F in Tab. 3), we observe improvements across most metrics (see comparisons from Setup C to E and from Setup D to F).
| IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | ||
| Channel-wise input | 68.7 | 58.5 | 33.4 | 0.7348 | 0.2959 |
| Channel-wise feature | 72.6 | 58.4 | 27.9 | 0.7372 | 0.2598 |
| Embedding feature | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
Where to inject heatmap
We explore various methods for incorporating the heatmap into the model and present the results in Tab. 4. First, we examine the channel-wise input approach, which involves concatenating the heatmap with the input image and feeding a 4-channel input into the image encoder. However, this method leads to a slight drop in performance. This is likely because the image encoder struggles to extract meaningful supervision from the fourth channel, as the heatmap is sparsely represented compared to the image content. In the second approach, channel-wise feature, we adopt the head prompting strategy proposed in [51]. Specifically, we use an embedding layer to learn heatmap embeddings, which are then added to the visual tokens. While this method is computationally efficient and does not require retraining the image encoder, the resulting performance improvement for our task is minimal. Finally, in the embedding feature setup, we introduce an additional CNN-based encoder to process the heatmap. The resulting heatmap embedding is passed through a projection head and then concatenated with the image and text embeddings before being fed into the transformer encoder. This approach yields the most substantial performance improvement among the methods evaluated.
| Ensemble method | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Confidence-Only | 73.4 | 64.1 | 34.1 | 0.7735 | 0.2470 |
| CLIP-Only (Top-1) | 73.5 | 64.2 | 35.4 | 0.7932 | 0.2391 |
| CLIP-Only (Top-2 + Threshold=0.95) | 73.8 | 64.1 | 35.5 | 0.7938 | 0.2377 |
| CLIP Fusion | 73.6 | 64.4 | 34.5 | 0.7379 | 0.2422 |
| CAPE | 75.0 | 65.4 | 35.7 | 0.7983 | 0.2476 |
CLIP-Aware Pointing Ensemble (CAPE)
We evaluate several ensemble strategies to determine the final referent bounding box from the outputs of the two pointing models and present the results in Tab. 5. (1) Confidence-Only: selects the box with the highest model confidence score among Top-1 prediction of both models. (2) CLIP-Only (Top-1): chooses the box with the highest CLIP similarity score between the text and Top-1 predictions from both models. (3) CLIP-Only (Top-2 + Threshold): extends the previous method by also considering the top-2 predictions from each model, provided their confidence scores exceed . (4) CLIP Fusion: computes a normalized CLIP score and sums it with the model’s confidence score. The box with the highest total score is selected. This method considers the top-2 predictions from both models. (5) CAPE: a hybrid method that combines approaches (3) and (4). For small objects, defined by [11] as occupying less than 0.48% of the image area, CAPE uses method (4), since CLIP tends to perform unreliably on very small referents. For all other objects, it uses method (3). This adaptive strategy balances the strengths of both scoring signals across varying object sizes. Please note that the threshold used in method (3), as well as the decision to apply method (4) for small objects, were determined based on performance analysis on the validation set to avoid overfitting to the test set.
Finetuned vs. frozen text encoder
We investigate the impact of freezing the text encoder during training. While freezing the text encoder reduces computational cost, mitigates overfitting in low-resource settings, and stabilizes the training (see Appendix D), we observe that finetuning consistently yields better performance across almost all evaluation metrics (see Tab. 6). This suggests that adapting the text features to the specific task of embodied reference understanding provides a meaningful benefit.
| Text Encoder | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Frozen, Setup A | 69.8 | 59.5 | 33.1 | 0.7363 | 0.2994 |
| Finetuned, Setup A | 71.2 | 60.1 | 32.8 | 0.7409 | 0.2662 |
| Frozen, Setup F | 71.3 | 62.1 | 33.7 | 0.7382 | 0.2634 |
| Finetuned, Setup F | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
How to create heatmap
We evaluate the Conic Attention Heatmap with and cone angles, as well as the Gaussian Ray Heatmap, which yields a more localized activation. Empirically, the Gaussian Ray Heatmap achieved the best performance (see Tab. 7). This suggests that broader attention regions may be less effective for the model, likely due to the high density of objects in typical scenes. Therefore, larger heatmaps tend to enclose multiple distractors, reducing the discriminative value of the spatial signal.
| Heatmap style | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
| Gaussian Ray Heatmap () | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
| Conic Attention Heatmap () | 70.0 | 60.0 | 33.9 | 0.7392 | 0.2632 |
| Conic Attention Heatmap () | 71.5 | 59.4 | 30.8 | 0.7389 | 0.2764 |
5 Conclusion
We tackled the challenge of Embodied Reference Understanding by addressing the limitations of existing single-line pointing assumptions. Our proposed dual-model framework explicitly leverages both head-to-fingertip and wrist-to-fingertip lines, represented as Gaussian ray heatmaps, to provide richer supervision beyond following pointing line and improve pointing target detection. By integrating these complementary signals through our CLIP-Aware Pointing Ensemble (CAPE) module and enhancing the model with an auxiliary object center prediction task, we achieve notable improvements on the YouRefIt benchmark. Our results demonstrate that combining multiple pointing cues and multimodal signals leads to more accurate and robust referent understanding in complex visual scenes.
Limitations
Due to the lack of explicit depth information, the model occasionally selects an incorrect object that lies along the pointing direction but is not the intended referent. We plan to address this depth-related issue in future work. Additionally, the use of a CNN-based image backbone leads to suboptimal precision in bounding box predictions. This limitation can be mitigated by adopting a more advanced backbone architecture, such as the Swin Transformer.
Supplementary Material
Appendix A Problem Analysis
How important is ERU task?
Embodied Reference Understanding (ERU) is a challenging multimodal task that combines visual, linguistic, and embodied cues (like gestures / pointing) to predict a target object in a scene. Despite recent progress, several key problems and open challenges remain. Natural language instructions are often vague or underspecified. Therefore, grounding models may detect multiple candidate objects or fail to resolve the correct referent altogether. Another problem is that when multiple instances of the same object class exist, models struggle to disambiguate. This results in over-detection (selecting all candidates) or under-confidence in selecting the target.
YouRefIt [11] introduced the ERU task, along with a multimodal dataset and a baseline approach to address it. However, the proposed method demonstrated limited performance, achieving a mean Average Precision (mAP) of only under an IoU threshold of . Subsequently, Touch-Line Transformer [35] proposed the concept of a Virtual Touch Line (VTL), based on the observation that the referent is typically aligned with the head-to-fingertip (H2F) line rather than the elbow-to-wrist (E2W) line. Since models struggle to follow pointing instructions, they propose using VTL prediction with a dedicated prediction head to encourage the model to focus more on the corresponding line. This approach led to a significant improvement in performance. However, our empirical analysis confirms that while the referent often aligns with the head-to-fingertip line, this is not always the case due to several factors:
If the person is looking elsewhere, such as at the camera, a robot, or another person, the alignment no longer holds.
The camera angle, consequently the image perspective, affects this alignment.
If the person is very close to the object or does not fully extend their arm while pointing the alignment may again be lost.
As a result, we find that in some cases, the referent aligns with the wrist-to-fingertip line. This indicates the need to incorporate both sources of information either in model training or decision step.
Appendix B Heatmap Generation
The heatmap generation method is a critical component of our approach. Recent work in this field [35] has shown that the referent often aligns with the head-to-fingertip line, referred to as the Virtual Touch-Line (VTL). Based on this insight, we generate a Gaussian Ray Heatmap using the head-to-fingertip line as a reference. While this assumption generally holds true, particularly when a person clearly points toward an object with an extended arm, it is not universally reliable. Variations occur both in this benchmark and in real-world scenarios. For instance, the alignment may be disrupted if the person is looking elsewhere (e.g., at a robot, camera, or another person). Similarly, camera perspective can also distort the alignment. In some cases, we observe that the wrist-to-fingertip line aligns more accurately with the referent than the head-to-fingertip line. To address this variability, we leverage both types of information and generate separate heatmaps for each sample accordingly.
: Heatmap from head-to-fingertip.
: Heatmap from wrist-to-fingertip.
B.1 Head-to-Fingertip Heatmap
To generate this heatmap, we use the head and fingertip as reference points to draw a line starting from the head and extending toward the edge of the image in the direction of the fingertip. In the YouRefIt dataset [11], we have access to detailed annotations from [11, 35], which include coordinates for the eyes, elbows, wrists, and fingertips. We utilize these annotations to construct the reference line and generate the corresponding heatmap. In real-world applications where such annotations are not available, estimating these key points is straightforward. For the eye position, face detection followed by selecting the center point of the detected face yields comparable performance. Likewise, for wrist and fingertip locations, modern pose estimation models provide reliable predictions.
B.2 Wrist-to-Fingertip Heatmap
We follow the same pipeline used for generating the head-to-fingertip heatmap, but replace the eye coordinates with wrist coordinates and the fingertip coordinates remain the same.
B.3 Merging
For the unimodal experiments, we merge the head-to-fingertip and wrist-to-fingertip heatmaps to combine information from different pointing directions. However, this approach yields lower accuracy compared to using two separate models. We hypothesize that this is primarily because the model cannot inherently determine which pointing line is more appropriate for each sample.
B.4 Heatmap Injection Methods
We investigate three different methods for injecting heatmaps and conduct experiments using the head-to-fingertip heatmap.
Channel-wise input
In this method, we concatenate the input image with the heatmap and feed it into the image encoder, resulting in a 4-channel input. Consequently, this approach is not compatible with pretrained image encoders unless they are finetuned, as it requires adapting the model to handle the additional channel. Experimental results show that this method yields the least accurate performance.
Channel-wise feature
In this method, we follow a recently proposed approach [51] that has shown strong performance in the gaze target estimation literature. Specifically, we use a single learnable embedding layer to obtain embeddings of the heatmap, which are then added to the image embeddings.
Embedding feature
In this approach, we propose using an additional encoder. Given the sparse nature of the heatmap representation, we select a simpler encoder than the main image encoder. In this work, we employ ResNet-18 as the heatmap encoder. After obtaining the embedded representation, we pass it through a projection layer to match the required dimensionality. We then concatenate it with the image and text embeddings and feed the combined representation into the transformer encoder.
Appendix C Ensemble Methods
In this section, we present the details of the ensemble methods we explored. Each of our models produces predictions, with each prediction accompanied by a confidence score. Before performing ensembling, we sort these predictions in descending order based on their confidence scores. We denote the sorted prediction list from model as , and from model as . Similarly, we denote the confidence score of model as , and from model as .
C.1 Confidence-Only
In the Confidence-Only approach, we focus solely on the top-1 prediction from each model. We take the bounding box with the highest confidence score from and another from . These two predictions are then compared, and the one with the higher confidence score is selected as the final prediction.
C.2 CLIP-Only (Top-1)
In this method, similar to the Confidence-Only approach, we select the bounding box with the highest confidence score from each model. Next, we compute the CLIP similarity score between each predicted bounding box and the input text. The bounding box with the higher CLIP similarity score is then chosen as the final prediction.
where is the embedded representation of cropped predicted bounding box of with the highest score and is the embedded representation of the text input.
where is the embedded representation of cropped predicted bounding box of with the highest score.
C.3 CLIP-Only (Top-2 + Threshold)
This method is similar to CLIP-Only (Top-1), where we take predictions from both models and compute the CLIP similarity score between the cropped bounding boxes and the input text. However, unlike the Top-1 variant, we also consider the Top-2 predictions from each model. However, we do not always include the second predictions; instead, we apply a confidence threshold determined empirically on the validation set, . If a model’s second-highest prediction has a confidence score greater than or equal to this threshold, it is included in the comparison. Finally, we select the bounding box with the highest CLIP similarity score as the final prediction. The algorithm for this method is presented in Algorithm C.3.
C.4 CLIP Fusion
In this method, we leverage both the CLIP similarity score and the model’s confidence score to make the final prediction. Specifically, we compute a hybrid score for each bounding box by summing its confidence score and its CLIP similarity score. We use the Top-2 predictions from both models for this process. However, the CLIP similarity scores and model confidence scores operate on different scales: while confidence scores are normalized in the range , CLIP similarity scores typically range around . To balance their influence and prevent the CLIP score from dominating, we scale the CLIP similarity scores by a factor of . The scaled CLIP score is then added to the corresponding confidence score to compute the final hybrid score for each bounding box. Finally, we select the bounding box with the highest hybrid score as the final prediction. The algorithm for this method is presented in Algorithm C.4.
C.5 CLIP-Aware Pointing Ensemble (CAPE)
In this method, we combine the CLIP-Only (Top-2 + Threshold) and CLIP Fusion approaches. Through empirical analysis on validation set, we observe that for small objects, the CLIP Fusion method yields more accurate results than other ensemble strategies. This is primarily because CLIP tends to perform less reliably on smaller objects. Therefore, we apply CLIP Fusion to not rely on only CLIP-score but also consider the confidence score. To determine whether an object is small, we follow the definition provided by [11]: if the object occupies less than of the total image area, we classify it as small and use the CLIP Fusion ensemble. For all other cases, we apply the CLIP-Only (Top-2 + Threshold) method. The algorithm is presented in Algorithm C.5.
| Name | Value |
| Learning rate vision backbone | |
| Learning rate - text backbone | |
| Learning rate | 1e-5 |
| 2 | |
| 1 | |
| 10 | |
| 10 | |
| 1 | |
| 1 | |
| Transformer encoder layer | 6 |
| Transformer decoder layer | 6 |
| Number of attention heads per layer | 8 |
| MLP dimension | 2048 |
| Dropout | p = 0.1 |
| Weight decay | 1e-4 |
| Batch size | 4 |
| Text encoder | Roberta-Base |
| Image encoder | ResNet-101 |
| Heatmap encoder | ResNet-18 |
| Position embedding | Sine |
| Input dim of transformer encoder | 256 |
| Number of queries | 20 |
Appendix D Finetuning or Frozen Text Encoder?
Based on multiple training experiments, we observed that finetuning the text encoder sometimes leads to instability, represented by the pink line in Fig. E.5. In contrast, freezing the text encoder results in consistently stable training, as shown by the green and orange lines in Fig. E.5. Therefore, achieving stable training with a finetuned text encoder requires several attempts. However, despite this challenge, finetuning the text encoder ultimately yields better performance.
Appendix E Hyperparameters and Other Training Parameters
In Tab. C.8, we present empirically determined hyperparameters and other key training details to support reproducibility.
Appendix F More Results
In Tab. F.9, we report the mAP at three different thresholds for all ablation setups, categorized by object size. Similarly, in Tab. F.10, we present the CLIP scores and metrics for the same setups, also with respect to object size.
In Fig. F.6, we present examples of failure cases. In the first row, although our model correctly detects the target object, CAPE selects the prediction from , which is incorrect in this case. In the second row, the bottles are extremely difficult to distinguish, even with explicit pointing cues. Nevertheless, the predicted bounding boxes cover other bottles adjacent to the GT one, which is acceptable, as they overlap with the pointing line and are semantically consistent with the text. In the third row, both of our models, as well as the Touch-Line models, produce incorrect predictions. However, guided by the pointing information, the predictions from and lie on the pointing line and are very close to the GT object. In contrast, the Touch-Line predictions are entirely off the pointing line and significantly distant from the GT object.
| IoU Threshold for mAP | 0.25 | 0.50 | 0.75 | ||||||||||||
| Setup | Object Sizes | Heatmap Injection | Ensemble | All | S | M | L | All | S | M | L | All | S | M | L |
| Touch-Line-EWL [35] | - | - | 69.5 | 56.6 | 71.7 | 80.0 | 60.7 | 44.4 | 66.2 | 71.2 | 35.5 | 11.8 | 38.9 | 55.0 | |
| Touch-Line-VTL [35] | - | - | 71.1 | 55.9 | 75.5 | 81.7 | 63.5 | 47.0 | 70.2 | 73.1 | 39.0 | 13.4 | 45.2 | 57.8 | |
| 1 | Baseline | - | - | 71.2 | 59.8 | 73.0 | 80.9 | 60.1 | 43.2 | 66.4 | 70.5 | 32.8 | 8.6 | 35.4 | 53.4 |
| 2 | Setup 1 + object center prediction | - | - | 70.8 | 59.5 | 73.2 | 79.5 | 61.7 | 45.8 | 67.4 | 71.5 | 34.6 | 14.0 | 38.1 | 51.1 |
| 3 | Setup 1 + W2F heatmap | Embedding feature | - | 68.9 | 55.5 | 74.0 | 77.4 | 59.7 | 42.0 | 68.1 | 69.3 | 32.4 | 8.7 | 37.8 | 50.6 |
| 4 | Setup 1 + H2F heatmap | Embedding feature | - | 71.9 | 57.5 | 74.1 | 84.1 | 62.8 | 43.3 | 70.1 | 75.1 | 33.8 | 10.0 | 35.6 | 55.3 |
| 5 | Setup 2 + W2F heatmap | Embedding feature | - | 69.6 | 57.7 | 74.9 | 76.4 | 60.7 | 45.5 | 69.0 | 68.2 | 31.5 | 12.1 | 35.3 | 47.0 |
| 6 | Setup 2 + H2F heatmap | Embedding feature | - | 72.9 | 59.5 | 77.8 | 81.4 | 62.3 | 44.3 | 70.9 | 42.0 | 35.1 | 10.4 | 39.3 | 55.3 |
| 7 | Setup 1 + W2F heatmap + H2F heatmap | Embedding feature | - | 70.2 | 59.0 | 71.4 | 80.2 | 60.2 | 44.5 | 65.6 | 70.5 | 33.8 | 11.9 | 39.3 | 50.0 |
| 8 | Setup 6 | Channel-wise input | - | 68.7 | 58.5 | 71.9 | 75.6 | 58.5 | 44.8 | 65.0 | 65.9 | 33.4 | 13.9 | 37.4 | 48.6 |
| 9 | Setup 6 | Channel-wise feature | - | 72.6 | 60.0 | 77.5 | 80.4 | 58.4 | 38.8 | 68.7 | 68.1 | 27.9 | 8.5 | 32.6 | 42.5 |
| 10 | Setup 5 + Setup 6 | Embedding feature | Confidence-Only | 73.4 | 62.7 | 78.6 | 79.0 | 64.1 | 49.0 | 72.7 | 71.0 | 34.1 | 13.4 | 38.2 | 50.5 |
| 11 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-1) | 73.5 | 59.7 | 79.9 | 81.2 | 64.2 | 46.5 | 74.3 | 72.2 | 35.4 | 12.7 | 39.8 | 53.4 |
| 12 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-2 + Threshold) | 73.8 | 59.7 | 80.2 | 81.6 | 64.1 | 46.3 | 74.3 | 72.2 | 35.5 | 12.4 | 40.1 | 53.9 |
| 13 | Setup 5 + Setup 6 | Embedding feature | CLIP Fusion | 73.6 | 63.2 | 78.4 | 79.6 | 64.4 | 49.3 | 72.8 | 71.7 | 34.5 | 13.4 | 38.5 | 51.7 |
| 14 | Setup 5 + Setup 6 | Embedding feature | CAPE | 75.0 | 63.2 | 80.2 | 81.8 | 65.4 | 49.5 | 74.3 | 72.7 | 35.7 | 13.4 | 40.1 | 53.4 |
| 15 | Setup 6 w/ frozen text encoder | Embedding feature | - | 71.3 | 56.0 | 74.9 | 83.1 | 62.1 | 42.5 | 69.5 | 74.4 | 33.7 | 10.2 | 35.8 | 54.6 |
| 16 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 70.0 | 54.2 | 72.2 | 83.3 | 60.0 | 42.0 | 66.6 | 71.5 | 33.9 | 12.2 | 40.6 | 48.8 |
| 17 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 71.5 | 58.7 | 73.5 | 82.1 | 59.4 | 41.8 | 64.7 | 71.7 | 30.8 | 7.0 | 32.4 | 52.7 |
| Ours - Final | Embedding feature | CAPE | 75.0 | 63.2 | 80.2 | 81.8 | 65.4 | 49.5 | 74.3 | 72.7 | 35.7 | 13.4 | 40.1 | 53.4 | |
| CLIP Score | |||||||||||
| Setup | Object Sizes | Heatmap Injection | Ensemble | All | S | M | L | All | S | M | L |
| Touch-Line-EWL [35] | - | - | 0.7369 | 0.6936 | 0.7305 | 0.7845 | 0.3168 | 0.3006 | 0.2903 | 0.3564 | |
| Touch-Line-VTL [35] | - | - | 0.7368 | 0.6925 | 0.7321 | 0.7846 | 0.2843 | 0.2809 | 0.2276 | 0.3393 | |
| 1 | Baseline | - | - | 0.7409 | 0.7003 | 0.7341 | 0.7861 | 0.2662 | 0.2182 | 0.2506 | 0.3266 |
| 2 | Setup 1 + object center prediction | - | - | 0.7339 | 0.6947 | 0.7277 | 0.7787 | 0.2707 | 0.2201 | 0.2692 | 0.3222 |
| 3 | Setup 1 + W2F heatmap | Embedding feature | - | 0.7353 | 0.6934 | 0.7297 | 0.7812 | 0.2984 | 0.2763 | 0.2616 | 0.3533 |
| 4 | Setup 1 + H2F heatmap | Embedding feature | - | 0.7375 | 0.6936 | 0.7333 | 0.7852 | 0.2694 | 0.2649 | 0.2233 | 0.3172 |
| 5 | Setup 2 + W2F heatmap | Embedding feature | - | 0.7344 | 0.6939 | 0.7306 | 0.7780 | 0.2770 | 0.2568 | 0.2319 | 0.3382 |
| 6 | Setup 2 + H2F heatmap | Embedding feature | - | 0.7373 | 0.6953 | 0.7343 | 0.7819 | 0.249 | 0.222 | 0.2202 | 0.3024 |
| 7 | Setup 1 + W2F heatmap + H2F heatmap | Embedding feature | - | 0.7345 | 0.6965 | 0.7277 | 0.7780 | 0.2744 | 0.2540 | 0.2314 | 0.3335 |
| 8 | Setup 6 | Channel-wise input | - | 0.7348 | 0.6933 | 0.7273 | 0.7828 | 0.2959 | 0.2437 | 0.2642 | 0.3767 |
| 9 | Setup 6 | Channel-wise feature | - | 0.7372 | 0.6953 | 0.7311 | 0.7839 | 0.2598 | 0.2289 | 0.2119 | 0.3339 |
| 10 | Setup 5 + Setup 6 | Embedding feature | Confidence-Only | 0.7735 | 0.7441 | 0.7843 | 0.7923 | 0.247 | 0.2176 | 0.2042 | 0.3141 |
| 11 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-1) | 0.7932 | 0.7688 | 0.8010 | 0.8098 | 0.2391 | 0.2242 | 0.2011 | 0.2878 |
| 12 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-2 + Threshold) | 0.7379 | 0.7967 | 0.7343 | 0.7819 | 0.2422 | 0.2158 | 0.2009 | 0.3060 |
| 13 | Setup 5 + Setup 6 | Embedding feature | CLIP Fusion | 0.7742 | 0.7927 | 0.7455 | 0.7842 | 0.2434 | 0.2137 | 0.2028 | 0.3090 |
| 14 | Setup 5 + Setup 6 (Final Model) | Embedding feature | CAPE | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
| 15 | Setup 6 w/ frozen text encoder | Embedding feature | - | 0.7382 | 0.6954 | 0.7323 | 0.7855 | 0.2634 | 0.2626 | 0.2233 | 0.3006 |
| 16 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.7387 | 0.6962 | 0.7325 | 0.7860 | 0.2695 | 0.2685 | 0.2344 | 0.3028 |
| 17 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.7391 | 0.6970 | 0.7332 | 0.7863 | 0.2765 | 0.2750 | 0.2499 | 0.3026 |
| Ours - Final | Embedding feature | CAPE | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 | |

### v4

@@
 Abstract
-We address the problem of Embodied Reference Understanding, which involves predicting the object that a person in the scene is referring to through both pointing gesture and language.
-Accurately identifying the referent requires multimodal understanding: integrating textual instructions, visual pointing, and scene context.
-However, existing methods often struggle to effectively leverage visual clues for disambiguation.
-We also observe that, while the referent is often aligned with the head-to-fingertip line, it occasionally aligns more closely with the wrist-to-fingertip line.
-Therefore, relying on a single line assumption can be overly simplistic and may lead to suboptimal performance.
+We address Embodied Reference Understanding, the task of predicting the object a person in the scene refers to through pointing gesture and language.
+This requires multimodal reasoning over text, visual pointing cues, and scene context, yet existing methods often fail to fully exploit visual disambiguation signals.
+We also observe that while the referent often aligns with the head-to-fingertip direction, in many cases it aligns more closely with the wrist-to-fingertip direction, making a single-line assumption overly limiting.
 To address this, we propose a dual-model framework, where one model learns from the head-to-fingertip direction and the other from the wrist-to-fingertip direction.
-We further introduce a Gaussian ray heatmap representation of these lines and use them as input to provide a strong supervisory signal that encourages the model to better attend to pointing cues.
-To combine the strengths of both models, we present the CLIP-Aware Pointing Ensemble module, which performs a hybrid ensemble based on CLIP features.
-Additionally, we propose an object center prediction head as an auxiliary task to further enhance referent localization.
-We validate our approach through extensive experiments and analysis on the benchmark YouRefIt dataset, achieving an improvement of approximately 4 mAP at the 0.25 IoU threshold.
+We introduce a Gaussian ray heatmap representation of these lines and use them as input to provide a strong supervisory signal that encourages the model to better attend to pointing cues.
+To fuse their complementary strengths, we present the CLIP-Aware Pointing Ensemble module, which performs a hybrid ensemble guided by CLIP features.
+We further incorporate an auxiliary object center prediction head to enhance referent localization.
+We validate our approach on YouRefIt, achieving 75.0 mAP at 0.25 IoU, alongside state-of-the-art CLIP and scores, and demonstrate its generality on unseen CAESAR and ISL Pointing, showing robust performance across benchmarks.
 1 Introduction
@@
 It plays a key role in real-world applications like human-robot interaction, assistive robotics, and augmented reality, where systems must understand which object a person refers to.
-While grounding models [37] have made significant progress in detecting objects mentioned in natural language, they often fall short in ERU task, particularly in ambiguous scenes.
+While grounding models [39] and LMMs [3, 4, 6, 56] have made significant progress in detecting objects mentioned in natural language, they often fall short in ERU task, particularly in ambiguous scenes.
 When multiple instances of the same object type are present, these models tend to detect all matching candidates without the ability to disambiguate and identify the specific target intended by the user.
 Moreover, when the textual instruction itself is vague or ambiguous, grounding models struggle even further, often failing to identify the correct target object or producing no confident prediction at all due to the inherent ambiguity.
-This limitation highlights the need for additional disambiguation cues, embodied gesture signals, that can help resolve referential ambiguity and enable accurate identification of the intended object.
-Therefore, both embodied gesture signals and language reference are crucial to identify the referent.
-In ERU, two main challenges arise.
-(1) identifying candidate objects from the text, as in standard grounding tasks, and (2) interpreting the pointing cue from visual input to guide the final prediction.
-The latter requires understanding human pose, inferring pointing direction, and handling visual complexities like perspective, occlusion, and depth.
-In [11], first time ERU task is defined and a multimodal YouRefIt dataset is proposed.
-They also proposed a method that combines textual and visual inputs along with PAF [7] and predicted saliency maps [33] to help localize the human and narrow down the potential referent.
-While these inputs lead to improved performance, the model still exhibits suboptimal accuracy in many cases, particularly due to its inability to properly follow the pointing instructions.
-Building on this, the Touch-Line Transformer [35] introduces a novel approach by predicting a Virtual Touch-Line and applying losses that encourage the model to focus on the head-to-fingertip direction, which they find often aligns closely with the intended referent [24, 44].
-Although this method achieves significant gains, it still faces challenges.
-The model occasionally overlooks the pointing line, as it is not explicitly provided as input.
-Instead, it predicts head, fingertips, and arm position to encourage attention in that direction.
-More importantly, the head-to-fingertip direction does not always align with the target.
-In some scenarios, the accurate pointing direction follows a line from the hand alone (e.g., wrist-to-fingertip), rather than from the head (Fig.
-1).
-These observations suggest that while head-to-fingertip direction offers valuable guidance, it is not sufficient on its own to fully resolve the complexity of the task.
+This limitation highlights the need for additional disambiguation cues that can help resolve referential ambiguity and enable accurate identification of the intended object.
+Therefore, both embodied gesture signals (pointing) and language reference are crucial to identify the referent, aligning with early multimodal interaction studies [72, 61].
+The ERU task was first defined in [13] with the YouRefIt dataset, and a method combining textual and visual inputs with PAF [8] and predicted saliency maps [36] was proposed.
+While effective, the model often fails to follow pointing instructions accurately.
+The Touch-Line Transformer [38] addresses this by predicting a Virtual Touch-Line along the head-to-fingertip direction.
+However, it still faces limitations as the pointing line is not explicitly provided as input, despite its success [84], and the head-to-fingertip direction does not always align with the target.
+In some cases, accurate pointing is better represented from the hand alone (e.g., wrist-to-fingertip, Fig.
+1), highlighting the need for more reliable pointing guidance.
 To tackle these challenges, we propose a dual-model framework that leverages complementary pointing cues.
@@
 One model utilizes a head-to-fingertip heatmap, while the other uses a wrist-to-fingertip heatmap, capturing different aspects of the pointing behavior.
-Their predictions are integrated via a CLIP-Aware Pointing Ensemble (CAPE) module, which effectively combines the strengths of both models.
-Additionally, we introduce an auxiliary objective to predict the object center, providing an extra supervisory signal that further enhances grounding accuracy.
-Our contributions are: (1) We propose using heatmap as extra modality to guide our end-to-end reference object detection model to focus on pointing information more.
+Since these two models provide complementary information depending on the scenario, we introduce the CLIP-Aware Pointing Ensemble (CAPE) module to effectively combine the strengths of both models.
+In this module, we leverage the CLIP model to compute semantic similarity scores between the input text and candidate object images.
+We use CLIP because it was trained on a large-scale image-text dataset, learning a joint embedding space in which semantically matching images and text are closely aligned.
+Additionally, we introduce an auxiliary object center prediction head, providing a supervisory signal that guides the model to more accurately localize the pointed object.
+Our contributions are: (1) We propose using a pointing heatmap as an additional modality to guide our model to focus more effectively on pointing cues.
+Specifically, we use detected head, wrist, and fingertip points to construct a Gaussian Ray Heatmap that highlights the approximate pointing area.
+The embedding of this pointing heatmap is then extracted via a heatmap encoder and provided to the model.
 (2) We propose two parallel complementary models to focus on head-to-fingertip and wrist-to-fingertip pointing lines, to individually alleviate inherent challenges of the task.
-(3) We present CAPE module to effectively ensemble two complementary models.
-(4) We introduce object center prediction head to enhance prediction accuracy by focusing on finding the object without considering the precise bounding box.
+(3) We present CAPE module to effectively ensemble two complementary models as ensembling has shown strong performance in other fields [19, 82, 27, 78, 83].
+(4) We introduce an object center prediction head to improve prediction accuracy by emphasizing object localization without requiring precise bounding boxes.
 2 Related Work
@@
 Nonverbal Communication for Referent
-Some studies [55, 56, 57, 67, 6, 66, 25, 45, 61, 29, 53, 64, 63, 20, 60, 68, 58, 62, 59, 21, 51] use gaze target detection as a nonverbal cue, however, gaze alone is unreliable due to distractions and the absence of a clear pointing moment.
+Some studies [57, 58, 59, 60, 71, 7, 70, 28, 46, 65, 32, 54, 68, 67, 23, 64, 73, 62, 66, 63, 24, 52] use gaze target detection as a nonverbal cue, however, gaze alone is unreliable due to distractions and the absence of a clear pointing moment.
 It is typically localized frame by frame and often used to support conversation [5, 10].
@@
 [16] introduced an interactive robot dialogue system that uses multimodal interaction and pointing line estimation to accurately identify referent with an iterative correction process.
-Their follow-up work [15] extended this by recognizing unseen objects using a Region Proposal Network and VL-T5 multimodal network [13], moving beyond general object classes from benchmarks like MSCOCO [36] to attribute-based object recognition.
+Their follow-up work [17] extended this by recognizing unseen objects using a Region Proposal Network and VL-T5 multimodal network [15], moving beyond general object classes.
 On the other hand, several datasets have been collected to incorporate nonverbal cues, but most remain in simulated domains due to accessibility and controllability [29, 30, 2, 31].
@@
 3 Methodology
-Problem Definition In ERU, given an RGB image and a text input , we predict a bounding box corresponding to an object referenced by the text and indicated by a pointing gesture in the image.
+In ERU, given an RGB image and a text input , we predict a bounding box corresponding to an object referenced by the text and indicated by a pointing gesture in the image.
 We propose using an additional pointing heatmap to supervise the model better focus on the visual pointing reference.
@@
 Fig.
-2 illustrates our architecture, which consists of a large-scale text encoder, image encoder, and heatmap encoder.
+2 illustrates our overall approach, which consists of two parallel models ( and ; only one is shown since the sole difference is the heatmap encoder input).
+After processing the text, image, and heatmap inputs, the outputs of both models are passed to the CLIP-Aware Pointing Ensemble (CAPE) module for the final decision.
+CAPE combines the outputs of these two complementary models to enhance overall performance.
+Specifically, and each comprise a large-scale text encoder, image encoder, and heatmap encoder.
 The outputs of these encoders are concatenated and fed into a transformer [69] encoder block.
-Subsequently, a transformer decoder processes the encoder output to generate predictions through task-specific heads.
-In this work, we introduce two networks that share the same architecture but are trained independently.
-The first network () receives a heatmap generated from the wrist-to-fingertip line, while the second () uses a heatmap derived from the head-to-fingertip line.
-Finally, a CLIP-Aware Pointing Ensemble (CAPE) module combines the outputs of these two complementary models to enhance overall performance.
+A subsequent transformer decoder then processes the encoder output to generate predictions via task-specific heads.
 Encoders
@@
 Since the pointing heatmap () is a sparse representation, we choose a lightweight encoder, ResNet-18 [25], to embed it: This choice not only improves the computational efficiency of the network but also helps prevent redundant information in the final concatenated feature representation, which could otherwise harm the model’s learning capacity.
-Transformer encoder
+Transformer encoder and decoder
 After obtaining the embeddings from three encoders, we first apply a convolution to each embedding individually to project their channel dimensions to .
 Next, we flatten the spatial dimensions of the embedding into a single dimension, converting them into sequences of tokens.
-For example, for the heatmap embedding: .
+For example, for the heatmap embedding: , where is flatten and is convolution.
 We then concatenate the three embeddings at the sequence level to form the final representation, which is fed into the transformer encoder.
 It processes this concatenated input to learn multimodal representations.
-Transformer decoder
 We feed the multimodal representation output from the transformer encoder into the transformer decoder.
@@
 This is particularly beneficial in real-world, cluttered scenes where gesture interpretation is inherently ambiguous.
-How should we create heatmap?
-The VTL [35] demonstrates superior performance in most cases, as it effectively captures the intended pointing direction by leveraging both the fingertip position and the head orientation.
-Human pointing gestures are typically guided by visual attention: we often align our finger with the target object based on our eye-level perspective [35].
-Consequently, the head–to-fingertip line serves as a strong proxy for the true pointing intent, integrating both gaze and gesture cues.
-However, the VTL is not universally reliable.
-In certain scenarios, the performer may be looking elsewhere (for instance, toward a robot or camera) rather than at the intended object.
-Additionally, in cases involving extreme viewing angles or occluded head poses, the head orientation may become ambiguous or misleading.
-In such situations, the local orientation of the hand, as approximated by the wrist–to-fingertip line, can offer complementary directional cues that help disambiguate the pointing gesture.
-We hypothesize that heatmaps derived from the head-to-fingertip and wrist-to-fingertip lines provide complementary information.
-Therefore, we generate heatmaps that incorporate both information.
-For the head-to-fingertip heatmap (), we use the eye coordinates as the starting point and draw a line toward the edge of the image, passing through the fingertip.
-Similarly, for the wrist-to-fingertip heatmap (), we start from the wrist and draw a line to the edge of the image through the fingertip point.
-To generate these heatmaps, we apply a Gaussian Ray Heatmap (See Appendix B).
+How should we create heatmaps?
+While prior work such as VTL [38] shows that head–to–fingertip cues can often capture the intended pointing direction, our goal is to design a heatmap representation that remains reliable across a wider range of real-world conditions.
+Human pointing gestures are guided by visual attention, and the head–to–fingertip line naturally encodes this gaze–gesture alignment.
+However, relying on this cue alone can be problematic when the performer is looking elsewhere (e.g., toward a robot or camera) or when the head pose becomes ambiguous due to extreme angles or occlusions.
+Moreover, close proximity to the object or incomplete arm extension can also break alignment.
+To address these limitations, our approach incorporates both the head–to–fingertip and wrist–to–fingertip lines.
+The wrist–to–fingertip line provides a complementary, locally grounded cue that remains informative even when head orientation is unreliable.
+By combining these two directional sources, we aim to obtain a more robust and consistent representation of pointing intent.
+Based on this hypothesis, we generate Gaussian Ray Heatmaps for both cues rather than relying solely on explicit line overlap.
+Both heatmaps extend a ray toward the image boundary through the fingertip, originating at the eyes for and at the wrist for .
+These heatmaps serve as complementary spatial supervision signals generated through our heatmap process (see Appendix A), capturing not only overlap with a single pointing line but also a broader region that reflects the performer’s focus of attention.
 How should we integrate these heatmaps?
-The next important question is how to integrate these heatmaps effectively.
-We explore two different approaches.
-In the first approach, we merge the two heatmaps () and feed the combined input into a single network to examine whether the model can implicitly learn to utilize both sources of information and determine the most appropriate direction based on the human pose.
-In the second approach, we train two identical networks independently and ensemble them: one using the head-to-finger heatmap () and the other using the wrist-to-fingertip heatmap ().
-We empirically find that the second approach yields better performance.
-As expected, it allows us to train two complementary networks.
-Notably, the network trained with the head-to-fingertip heatmap () demonstrates significant improvements.
-In contrast, the first approach results in only subtle enhancement, and the model fails to effectively leverage both sources of information to determine the pointing direction (see Tab.
-3 Setup G).
-Its behavior closely resembles the head-to-fingertip-only scenario.
-However, due to the presence of additional wrist-to-fingertip line in the merged heatmap, the model experiences contradictory information, ultimately leading to reduced accuracy than head-to-fingertip heatmap-only setup.
-Where should we inject these heatmaps?
-We empirically find that the most effective way to leverage heatmaps is to use a separate encoder to extract heatmap embeddings, which are then concatenated with the output of the image encoder.
-It enables us to preserve the pretrained image encoder and allows the dedicated heatmap encoder to extract rich, modality-specific information.
+We investigate two strategies for integrating the two heatmaps.
+The first merges them directly () and feeds the combined input into a single network, testing whether the model can implicitly resolve the two cues and infer the correct pointing direction from pose context.
+The second trains two identical networks separately, one with the head-to-fingertip heatmap () and one with the wrist-to-fingertip heatmap (), and ensembles their predictions.
+Empirically, the second strategy works substantially better.
+It yields two complementary models, with providing the strongest gains.
+In contrast, the merged-heatmap approach offers only marginal improvement.
+The model behaves similarly to the head-only variant, and the extra wrist-to-fingertip signal introduces conflicting directional cues that ultimately reduce accuracy compared to using the head-to-fingertip heatmap alone (see Tab.
+4, Setup G).
+We find that the most effective way to leverage heatmaps is to process them with a dedicated encoder whose embeddings are concatenated with the image encoder’s output.
+This design preserves the pretrained visual backbone while allowing the heatmap encoder to capture richer, modality-specific spatial cues.
 3.3 CLIP-Aware Pointing Ensemble (CAPE)
-We introduce an inference time ensemble module to increase the performance by benefiting from two complementary models, and .
+We introduce an inference time ensemble module to increase the performance by benefiting from two complementary models, and , since we find that they demonstrate strength in different scenarios.
 CAPE is an adaptive scoring method that combines model confidence and CLIP-based similarity in a size-aware manner effectively.
+We choose the CLIP model because it was trained to understand and evaluate image-text semantic similarity, which aligns well with our task.
 (1) For each prediction (top-2 predictions from both models), we compute a normalized CLIP score [49].
-We sum pointing models’ confidece scores with the CLIP scores to obtain the fused score (CLIP Fusion in Tab.
-5).
+We sum pointing models’ confidence scores with the CLIP scores to obtain the fused score (CLIP Fusion in Tab.
+6) since the models’ confidence scores also provide reliable insights.
 (2) We calculate CLIP scores for the top-1 predictions of both models.
@@
 Finally, we select the prediction with the highest CLIP score from among these candidate boxes (CLIP-Only Top-2 + Threshold=0.95 in Tab.
-5).
+6).
 CAPE applies strategy (1) when the referent is a small object (defined as occupying less than 0.48% of the image area, following [13]) to avoid relying solely on the CLIP prediction, as CLIP becomes less reliable for smaller objects.
 For all other objects, strategy (2) is used (see Sec.
-C.5 for details).
+B.5 for details).
 This hybrid approach leverages the strengths of both signals while adapting to object scale.
@@
 () are the center of the predicted bounding box.
+is cosine similarity.
 We compute the same similarity using GT bounding box to obtain .
@@
 and are soft token loss and contrastive loss respectively to help visual and textural signals alignment as in [33].
-We empirically determine the optimal coefficients on the validation set (See Appendix E).
+We empirically determine the optimal coefficients on the validation set (See Appendix D).
 | IoU Threshold for mAP | 0.25 | 0.50 | 0.75 | |||||||||
 | Object Sizes | All | S | M | L | All | S | M | L | All | S | M | L |
+| PaliGemma2 [56] | 58.8 | 29.0 | 53.5 | 75.8 | 46.9 | 22.1 | 50.8 | 68.0 | 31.7 | 6.2 | 34.1 | 54.8 |
+| Qwen2.5vl [4] | 38.9 | 17.0 | 41.8 | 58.0 | 31.0 | 11.1 | 33.6 | 48.1 | 20.0 | 5.7 | 19.8 | 34.5 |
 | Grounding DINO [39] | 57.9 | 38.0 | 60.9 | 74.9 | 54.9 | 35.7 | 59.3 | 69.6 | 42.3 | 22.7 | 45.9 | 58.4 |
@@
 Dataset
-For both training and testing, we use the image subset of the YouRefIt dataset [11], as the video portion is not publicly available.
-The dataset contains 2,950 training images and 1,245 test images.
-Annotations for bounding boxes, as well as eye, fingertip, elbow, and wrist coordinates, are provided by [11, 35].
+For both training and testing, we use the YouRefIt dataset [13], which contains 2,950 training and 1,245 test images.
+Text instructions and the annotations of bounding boxes, as well as pointing line coordinates, are provided by [13, 38].
+Additionally, we evaluate our approach on the unseen and more challenging ISL Pointing dataset [16], and a simulation dataset named CAESAR [29] (for results see appx.
+Tab.
+E.13) to assess its generalization capability.
 Evaluation
 For fair evaluation and comparison, we follow prior work [13] and report mean Average Precision (mAP) under three Intersection-over-Union (IoU) thresholds: 0.25, 0.50, and 0.75.
-A prediction is considered correct if its IoU with the GT exceeds the given threshold.
-Additionally, by following the literature, mAP scores are reported with respect to object size, categorized as Small (S), Medium (M), and Large (L).
-The object size categories are determined based on the ratio of the object area to the image area, using two thresholds [11]: and .
-An object is classified as small if its area ratio is below , medium if it falls between and , and large otherwise.
+Additionally, mAP scores are reported with respect to object size, categorized as Small (S), Medium (M), and Large (L) which based on the ratio of the object area to the image area using two thresholds: and .
 In addition to this standard evaluation protocol, we introduce two metrics for assessing referential grounding performance: CLIP score and center coordinate distance ().
@@
 We then calculate the cosine similarity between these feature vectors to quantify the semantic alignment between the predicted object and the text.
-For , we compute the L1 distance between the center coordinates of the predicted bounding box and those of the GT bounding box, providing a direct measure of spatial alignment.
+For , we compute the L1 distance between the center coordinates of the predicted and the GT bounding boxes, providing a direct measure of spatial alignment.
 Implementation
@@
 | Objet Sizes | All | S | M | L | All | S | M | L |
-| Touch-Line-EWL | 0.7369 | 0.6936 | 0.7305 | 0.7845 | 0.3168 | 0.3006 | 0.2903 | 0.3564 |
-| Touch-Line-VTL | 0.7368 | 0.6925 | 0.7321 | 0.7846 | 0.2843 | 0.2809 | 0.2276 | 0.3393 |
-| Ours | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
+| PaliGemma2 | 0.2488 | 0.2302 | 0.2446 | 0.2659 | 0.4684 | 0.5784 | 0.4307 | 0.4175 |
+| Qwen2.5vl-3b | 0.2408 | 0.2261 | 0.2388 | 0.2568 | 0.7288 | 0.7074 | 0.6893 | 0.7891 |
+| Grounding DINO | 0.2437 | 0.2267 | 0.2413 | 0.2621 | 0.3496 | 0.2946 | 0.4521 | 0.3104 |
+| Touch-Line-EWL | 0.2456 | 0.2312 | 0.2435 | 0.2615 | 0.3168 | 0.3006 | 0.2903 | 0.3564 |
+| Touch-Line-VTL | 0.2456 | 0.2308 | 0.2440 | 0.2615 | 0.2843 | 0.2809 | 0.2276 | 0.3393 |
+| Ours (CAPE) | 0.2661 | 0.2642 | 0.2671 | 0.2670 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
+| Method | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
+| PaliGemma2 | 47.2 | 39.5 | 31.6 | 0.2465 | 0.7449 |
+| Qwen2.5vl-3b | 32.5 | 32.1 | 29.2 | 0.2264 | 0.8418 |
+| Grounding DINO | 27.5 | 27.5 | 26.3 | 0.2082 | 0.7956 |
+| Touch-Line-EWL | 45.0 | 35.8 | 22.0 | 0.2436 | 0.5160 |
+| Touch-Line-VTL | 47.7 | 36.7 | 17.4 | 0.2473 | 0.5147 |
+| Ours, | 48.2 | 33.6 | 20.0 | 0.2453 | 0.4735 |
+| Ours, | 42.7 | 30.0 | 17.3 | 0.2449 | 0.6795 |
+| Ours, CAPE | 54.5 | 42.7 | 25.5 | 0.2642 | 0.4403 |
 4.1 Results
@@
 1 compares the performance of our method with existing approaches on the YouRefIt benchmark.
-Our model achieves SOTA results across all objects at IoU thresholds of 0.25 and 0.50, despite being trained for only 30 epochs on a single GPU with a batch size of 4.
-According to our experiments, this limited training setup affects bounding box precision, contributing to the suboptimal performance at , alongside limitations related to the backbone (see Grounding DINO in Tab.
-1 for backbone).
-In contrast, Touch-Line [35] was trained for 200 epochs with a batch size of 56.
+Our model achieves SOTA results in most cases, despite being trained for only 30 epochs on a single GPU with a batch size of 4.
+In contrast, Touch-Line models [38] were trained for 200 epochs with a batch size of 56.
 This clearly demonstrates the efficiency and effectiveness of our approach.
+However, our optimized training setup affects bounding box precision, contributing to the suboptimal performance at , alongside limitations related to the simpler backbone.
+At , Grounding DINO with a more complex backbone achieves the best performance.
+Further, we evaluate recent LMMs on this task, and they fall significantly short of our model’s performance, despite being among the largest models and trained for visual grounding purposes.
+These results show using complementary pointing heatmap modalities in addition to image and text is essential.
+Besides, despite having two separate models, our approach shows better inference-time performance (2.44 fps) compared to Paligemma2 (1.24fps) and Qwen2.5vl (0.5fps).
 Tab.
-2 presents a comparison between our method and the previous SOTA, Touch-Line, using two new metrics in this field: CLIP score and .
-The higher CLIP scores indicate that our predicted bounding boxes have stronger semantic alignment with the input text, suggesting better object identification and improved bounding box precision.
-Similarly, our method achieves the best performance under the metric, which quantifies the distance between the centers of the predicted and GT bounding boxes.
+2 presents further comparison using two new metrics in this field: CLIP score and .
+Higher CLIP scores indicate that our predicted bounding boxes have stronger semantic alignment with the input text, suggesting better object identification and improved bounding box precision.
+Similarly, our method achieves the best performance under the metric.
 This confirms that our model produces more spatially accurate predictions.
+In Tab.
+3, the experimental results on unseen ISL dataset show that our CAPE model surpasses the SOTA Touch-Line model as well as the evaluated LMMs, except for the IoU=0.75 setup.
+This demonstrates the strong generalization capability of our method.
 In Fig.
-3, we present samples images with predicted and GT bounding boxes.
-We compare our model with Touch-Line [35].
+3, we present sample images with predicted and GT bounding boxes.
+It is clearly seen from LMMs’ predictions, ambiguous text without pointing cues results in poor object localization outside of the pointing area and confusion with similar objects.
+For instance, Paligemma2 in third example and Qwen2.5vl in second and fourth examples detect wrong objects.
+On the other hand, the LMMs are strong at providing precise bounding box predictions due to extensive training and a robust backbone.
+Furthermore, even if Touch-Line model use pointing cues since it follows a pointing line instead of heatmaps, this let model to detect other objects in the pointing direction.
+In contrast, as shown in the examples, CAPE leverages explicit pointing heatmaps, enabling it to more accurately localize the intended referent.
+Also, it enables to choose final prediction from which of the more accurate model.
 | Setup | Method | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
-| A | Baseline | 71.2 | 60.1 | 32.8 | 0.7409 | 0.2662 |
-| B | A + object center prediction | 70.8 | 61.6 | 34.6 | 0.7339 | 0.2707 |
-| C | A + W2F heatmap | 68.9 | 59.7 | 32.4 | 0.7353 | 0.2984 |
-| D | A + H2F heatmap | 71.9 | 62.8 | 33.8 | 0.7375 | 0.2694 |
-| E | B + W2F heatmap () | 69.6 | 60.7 | 31.5 | 0.7344 | 0.2770 |
-| F | B + H2F heatmap () | 72.9 | 62.3 | 35.1 | 0.7373 | 0.2490 |
-| G | A + W2F heatmap + H2F heatmap | 70.2 | 60.2 | 33.8 | 0.7345 | 0.2744 |
-| H | Full model - Ensemble of E and F | 75.0 | 65.4 | 35.7 | 0.7983 | 0.2476 |
+| A | Baseline | 71.2 | 60.1 | 32.8 | 0.2469 | 0.2662 |
+| B | A + object center prediction | 70.8 | 61.6 | 34.6 | 0.2446 | 0.2707 |
+| C | A + W2F heatmap | 68.9 | 59.7 | 32.4 | 0.2451 | 0.2984 |
+| D | A + H2F heatmap | 71.9 | 62.8 | 33.8 | 0.2458 | 0.2694 |
+| E | B + W2F heatmap () | 69.6 | 60.7 | 31.5 | 0.2448 | 0.2770 |
+| F | B + H2F heatmap () | 72.9 | 62.3 | 35.1 | 0.2457 | 0.2490 |
+| G | A + W2F heatmap + H2F heatmap | 70.2 | 60.2 | 33.8 | 0.2448 | 0.2744 |
+| H | Full model - Ensemble of E and F | 75.0 | 65.4 | 35.7 | 0.2661 | 0.2476 |
 4.2 Analysis
 In this section, we provide a detailed analysis of the optimality of our design choices, including heatmap generation and the ensemble method.
-Specifically, we examine various heatmap injection strategies and ensemble approaches, among other factors.
 The baseline in Tab.
@@
 Object center prediction
-The object center prediction head and its associated loss yield improvements over the baseline in terms of mAP at and .
-However, they have a detrimental effect on mAP at , as well as on the CLIP Score and .
-Notably, when combined with the heatmap (Setup E and F in Tab.
-3), we observe improvements across most metrics (see comparisons from Setup C to E and from Setup D to F).
+The object center prediction improves mAP at and but reduces performance at , as well as CLIP Score and .
+Notably, when combined with heatmap (Setup E and F in Tab.
+4), it boosts most metrics.
+The object center prediction alone enhances geometric precision more than semantic alignment, while pairing it with spatially dense cues makes it a more effective refining signal.
 | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | ||
-| Channel-wise input | 68.7 | 58.5 | 33.4 | 0.7348 | 0.2959 |
-| Channel-wise feature | 72.6 | 58.4 | 27.9 | 0.7372 | 0.2598 |
-| Embedding feature | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
+| Channel-wise input | 68.7 | 58.5 | 33.4 | 0.2449 | 0.2959 |
+| Channel-wise feature | 72.6 | 58.4 | 27.9 | 0.2457 | 0.2598 |
+| Embedding feature | 72.9 | 62.2 | 35.1 | 0.2458 | 0.2490 |
 Where to inject heatmap
-We explore various methods for incorporating the heatmap into the model and present the results in Tab.
-4.
-First, we examine the channel-wise input approach, which involves concatenating the heatmap with the input image and feeding a 4-channel input into the image encoder.
-However, this method leads to a slight drop in performance.
-This is likely because the image encoder struggles to extract meaningful supervision from the fourth channel, as the heatmap is sparsely represented compared to the image content.
-In the second approach, channel-wise feature, we adopt the head prompting strategy proposed in [51].
-Specifically, we use an embedding layer to learn heatmap embeddings, which are then added to the visual tokens.
-While this method is computationally efficient and does not require retraining the image encoder, the resulting performance improvement for our task is minimal.
-Finally, in the embedding feature setup, we introduce an additional CNN-based encoder to process the heatmap.
-The resulting heatmap embedding is passed through a projection head and then concatenated with the image and text embeddings before being fed into the transformer encoder.
-This approach yields the most substantial performance improvement among the methods evaluated.
+We compare several heatmap injection strategies in Tab.
+5.
+The channel-wise input method, which feeds the heatmap as a fourth image channel, slightly hurts performance, likely because the heatmap is sparsely represented compared to the image content.
+Inspired by the head location prompting strategy in [52], the channel-wise feature approach adds learned heatmap embeddings to the visual tokens, which is lightweight but provides only limited gains.
+In contrast, the embedding feature setup uses a dedicated CNN to encode the heatmap before fusing it with image and text embeddings, and delivers the strongest improvements across all metrics.
+The experiments show that effective heatmap injection requires the model to interpret the heatmap as a structured spatial signal rather than as a raw image channel or a light additive bias on visual tokens.
+Besides, the improvements seen in the embedding feature method support the gains observed in the main ablations (Setups D and F), where the H2F heatmap consistently strengthens localization and reduces .
 | Ensemble method | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
-| Confidence-Only | 73.4 | 64.1 | 34.1 | 0.7735 | 0.2470 |
-| CLIP-Only (Top-1) | 73.5 | 64.2 | 35.4 | 0.7932 | 0.2391 |
-| CLIP-Only (Top-2 + Threshold=0.95) | 73.8 | 64.1 | 35.5 | 0.7938 | 0.2377 |
-| CLIP Fusion | 73.6 | 64.4 | 34.5 | 0.7379 | 0.2422 |
-| CAPE | 75.0 | 65.4 | 35.7 | 0.7983 | 0.2476 |
+| Confidence-Only | 73.4 | 64.1 | 34.1 | 0.2578 | 0.2470 |
+| CLIP-Only (Top-1) | 73.5 | 64.2 | 35.4 | 0.2644 | 0.2391 |
+| CLIP-Only (Top-2 + Threshold=0.95) | 73.8 | 64.1 | 35.5 | 0.2646 | 0.2377 |
+| CLIP Fusion | 73.6 | 64.4 | 34.5 | 0.2459 | 0.2422 |
+| CAPE | 75.0 | 65.4 | 35.7 | 0.2661 | 0.2476 |
 CLIP-Aware Pointing Ensemble (CAPE)
-We evaluate several ensemble strategies to determine the final referent bounding box from the outputs of the two pointing models and present the results in Tab.
-5.
-(1) Confidence-Only: selects the box with the highest model confidence score among Top-1 prediction of both models.
-(2) CLIP-Only (Top-1): chooses the box with the highest CLIP similarity score between the text and Top-1 predictions from both models.
-(3) CLIP-Only (Top-2 + Threshold): extends the previous method by also considering the top-2 predictions from each model, provided their confidence scores exceed .
-(4) CLIP Fusion: computes a normalized CLIP score and sums it with the model’s confidence score.
-The box with the highest total score is selected.
-This method considers the top-2 predictions from both models.
-(5) CAPE: a hybrid method that combines approaches (3) and (4).
-For small objects, defined by [11] as occupying less than 0.48% of the image area, CAPE uses method (4), since CLIP tends to perform unreliably on very small referents.
-For all other objects, it uses method (3).
-This adaptive strategy balances the strengths of both scoring signals across varying object sizes.
-Please note that the threshold used in method (3), as well as the decision to apply method (4) for small objects, were determined based on performance analysis on the validation set to avoid overfitting to the test set.
-Finetuned vs. frozen text encoder
-We investigate the impact of freezing the text encoder during training.
-While freezing the text encoder reduces computational cost, mitigates overfitting in low-resource settings, and stabilizes the training (see Appendix D), we observe that finetuning consistently yields better performance across almost all evaluation metrics (see Tab.
-6).
-This suggests that adapting the text features to the specific task of embodied reference understanding provides a meaningful benefit.
-| Text Encoder | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
-| Frozen, Setup A | 69.8 | 59.5 | 33.1 | 0.7363 | 0.2994 |
-| Finetuned, Setup A | 71.2 | 60.1 | 32.8 | 0.7409 | 0.2662 |
-| Frozen, Setup F | 71.3 | 62.1 | 33.7 | 0.7382 | 0.2634 |
-| Finetuned, Setup F | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
-How to create heatmap
-We evaluate the Conic Attention Heatmap with and cone angles, as well as the Gaussian Ray Heatmap, which yields a more localized activation.
-Empirically, the Gaussian Ray Heatmap achieved the best performance (see Tab.
-7).
-This suggests that broader attention regions may be less effective for the model, likely due to the high density of objects in typical scenes.
-Therefore, larger heatmaps tend to enclose multiple distractors, reducing the discriminative value of the spatial signal.
-| Heatmap style | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
-| Gaussian Ray Heatmap () | 72.9 | 62.2 | 35.1 | 0.7373 | 0.2490 |
-| Conic Attention Heatmap () | 70.0 | 60.0 | 33.9 | 0.7392 | 0.2632 |
-| Conic Attention Heatmap () | 71.5 | 59.4 | 30.8 | 0.7389 | 0.2764 |
+We evaluate several ensemble strategies for selecting the final referent box from the two pointing models, as shown in Tab.
+6.
+(1) Confidence-Only selects the Top-1 prediction with the highest model confidence.
+(2) CLIP-Only (Top-1) chooses the Top-1 box with the highest CLIP similarity.
+(3) CLIP-Only (Top-2 + Threshold) additionally considers each model’s top-2 predictions when their confidence scores exceed .
+(4) CLIP Fusion combines normalized CLIP and confidence scores from the top-2 predictions of both models and selects the box with the highest total score.
+(5) CAPE adaptively switches between (3) and (4): it uses (4) for small objects, where CLIP alone is unreliable, and (3) otherwise.
+The confidence threshold in (3) and the small-object rule in CAPE were tuned on the validation set to avoid test-time bias.
+Across these strategies, the hybrid design in CAPE aligns with observations from the main ablations: the two models provide complementary strengths, CLIP-based scoring refines semantic alignment for most objects, and confidence-based cues are especially important for small or ambiguous targets.
+This contributes to the strong gains seen in the full model (Setup H).
+Failures
+In Fig.
+4, we present examples of failure cases.
+In the first example (column), although our model correctly detects the target object, CAPE selects the prediction from , which is incorrect in this case.
+In the second example, both of our models, as well as the Touch-Line models, produce incorrect predictions.
+However, guided by the pointing information, the predictions from and lie on the pointing line and are very close to the GT object.
+In contrast, the Touch-Line predictions are entirely off the pointing line and significantly distant from the GT object.
+In the third column, the bottles are extremely difficult to distinguish, even with explicit pointing cues.
+Nevertheless, the predicted bounding boxes cover other bottles adjacent to the GT one, which is acceptable, as they overlap with the pointing line and are semantically consistent with the text.
 5 Conclusion
-We tackled the challenge of Embodied Reference Understanding by addressing the limitations of existing single-line pointing assumptions.
-Our proposed dual-model framework explicitly leverages both head-to-fingertip and wrist-to-fingertip lines, represented as Gaussian ray heatmaps, to provide richer supervision beyond following pointing line and improve pointing target detection.
-By integrating these complementary signals through our CLIP-Aware Pointing Ensemble (CAPE) module and enhancing the model with an auxiliary object center prediction task, we achieve notable improvements on the YouRefIt benchmark.
-Our results demonstrate that combining multiple pointing cues and multimodal signals leads to more accurate and robust referent understanding in complex visual scenes.
-Limitations
-Due to the lack of explicit depth information, the model occasionally selects an incorrect object that lies along the pointing direction but is not the intended referent.
-We plan to address this depth-related issue in future work.
-Additionally, the use of a CNN-based image backbone leads to suboptimal precision in bounding box predictions.
-This limitation can be mitigated by adopting a more advanced backbone architecture, such as the Swin Transformer.
+We address Embodied Reference Understanding by overcoming the limitations of relying on a single pointing line, using our dual-model framework to leverage complementary Gaussian ray heatmaps for richer supervision and improved target detection.
+Combined through the CLIP-Aware Pointing Ensemble (CAPE) and enhancing the model with an auxiliary object center prediction, our approach achieves strong improvements on the YouRefIt, ISL pointing, and CAESAR datasets.
+Our results demonstrate that combining multiple pointing cues and in general, multimodal signals, leads to more accurate and robust referent understanding in complex visual scenes.
+Beyond empirical gains, our work demonstrates a general principle of multimodal cue integration: fusing complementary spatial and semantic signals resolves ambiguities that single modalities cannot, providing a principle for robust referent understanding in complex visual scenes.
+Limitations and Future Work
+The model occasionally selects incorrect objects along the pointing direction due to the absence of depth information, and the CNN backbone limits bounding box precision.
+Future work could incorporate depth or 3D cues and adopt advanced backbones such as Swin Transformers.
+Additionally, formalizing a theoretical framework for multimodal cue integration and exploring adaptive weighting or attention mechanisms could further improve flexibility and robustness, allowing the system to optimally leverage multiple cues across diverse scenarios.
+Acknowledgment
+This work was supported in part by the European Union’s Horizon research and innovation program, project Meetween (101135798) and project DVPS (Diversibus Viis Plurima Solvo) (101213369), and KIT Campus Transfer GmbH (KCT) under contract from Carnegie-AI LLC.
 Supplementary Material
-Appendix A Problem Analysis
-How important is ERU task?
-Embodied Reference Understanding (ERU) is a challenging multimodal task that combines visual, linguistic, and embodied cues (like gestures / pointing) to predict a target object in a scene.
-Despite recent progress, several key problems and open challenges remain.
-Natural language instructions are often vague or underspecified.
-Therefore, grounding models may detect multiple candidate objects or fail to resolve the correct referent altogether.
-Another problem is that when multiple instances of the same object class exist, models struggle to disambiguate.
-This results in over-detection (selecting all candidates) or under-confidence in selecting the target.
-YouRefIt [11] introduced the ERU task, along with a multimodal dataset and a baseline approach to address it.
-However, the proposed method demonstrated limited performance, achieving a mean Average Precision (mAP) of only under an IoU threshold of .
-Subsequently, Touch-Line Transformer [35] proposed the concept of a Virtual Touch Line (VTL), based on the observation that the referent is typically aligned with the head-to-fingertip (H2F) line rather than the elbow-to-wrist (E2W) line.
-Since models struggle to follow pointing instructions, they propose using VTL prediction with a dedicated prediction head to encourage the model to focus more on the corresponding line.
-This approach led to a significant improvement in performance.
-However, our empirical analysis confirms that while the referent often aligns with the head-to-fingertip line, this is not always the case due to several factors:
-If the person is looking elsewhere, such as at the camera, a robot, or another person, the alignment no longer holds.
-The camera angle, consequently the image perspective, affects this alignment.
-If the person is very close to the object or does not fully extend their arm while pointing the alignment may again be lost.
-As a result, we find that in some cases, the referent aligns with the wrist-to-fingertip line.
-This indicates the need to incorporate both sources of information either in model training or decision step.
-Appendix B Heatmap Generation
-The heatmap generation method is a critical component of our approach.
-Recent work in this field [35] has shown that the referent often aligns with the head-to-fingertip line, referred to as the Virtual Touch-Line (VTL).
-Based on this insight, we generate a Gaussian Ray Heatmap using the head-to-fingertip line as a reference.
-While this assumption generally holds true, particularly when a person clearly points toward an object with an extended arm, it is not universally reliable.
-Variations occur both in this benchmark and in real-world scenarios.
-For instance, the alignment may be disrupted if the person is looking elsewhere (e.g., at a robot, camera, or another person).
-Similarly, camera perspective can also distort the alignment.
-In some cases, we observe that the wrist-to-fingertip line aligns more accurately with the referent than the head-to-fingertip line.
-To address this variability, we leverage both types of information and generate separate heatmaps for each sample accordingly.
+Appendix A Heatmap Generation
+The heatmap generation method is a key component of our approach.
+Prior work [38] shows that the referent often aligns with the head-to-fingertip line, which inspired our Gaussian Ray Heatmap generation.
+While this generally holds when a person points clearly with an extended arm, it is not always reliable.
+Alignment can be disrupted if the person looks elsewhere (e.g., at a robot, camera, or another person), due to camera perspective, or when a person is close to the object and points using wrist motion.
+In such cases, the wrist-to-fingertip line aligns more accurately with the referent.
+To handle this variability, we generate separate heatmaps for both cues.
 : Heatmap from head-to-fingertip.
 : Heatmap from wrist-to-fingertip.
-B.1 Head-to-Fingertip Heatmap
-To generate this heatmap, we use the head and fingertip as reference points to draw a line starting from the head and extending toward the edge of the image in the direction of the fingertip.
-In the YouRefIt dataset [11], we have access to detailed annotations from [11, 35], which include coordinates for the eyes, elbows, wrists, and fingertips.
-We utilize these annotations to construct the reference line and generate the corresponding heatmap.
-In real-world applications where such annotations are not available, estimating these key points is straightforward.
-For the eye position, face detection followed by selecting the center point of the detected face yields comparable performance.
-Likewise, for wrist and fingertip locations, modern pose estimation models provide reliable predictions.
-B.2 Wrist-to-Fingertip Heatmap
-We follow the same pipeline used for generating the head-to-fingertip heatmap, but replace the eye coordinates with wrist coordinates and the fingertip coordinates remain the same.
-B.3 Merging
-For the unimodal experiments, we merge the head-to-fingertip and wrist-to-fingertip heatmaps to combine information from different pointing directions.
-However, this approach yields lower accuracy compared to using two separate models.
-We hypothesize that this is primarily because the model cannot inherently determine which pointing line is more appropriate for each sample.
-B.4 Heatmap Injection Methods
-We investigate three different methods for injecting heatmaps and conduct experiments using the head-to-fingertip heatmap.
-Channel-wise input
-In this method, we concatenate the input image with the heatmap and feed it into the image encoder, resulting in a 4-channel input.
-Consequently, this approach is not compatible with pretrained image encoders unless they are finetuned, as it requires adapting the model to handle the additional channel.
-Experimental results show that this method yields the least accurate performance.
-Channel-wise feature
-In this method, we follow a recently proposed approach [51] that has shown strong performance in the gaze target estimation literature.
-Specifically, we use a single learnable embedding layer to obtain embeddings of the heatmap, which are then added to the image embeddings.
-Embedding feature
-In this approach, we propose using an additional encoder.
-Given the sparse nature of the heatmap representation, we select a simpler encoder than the main image encoder.
-In this work, we employ ResNet-18 as the heatmap encoder.
-After obtaining the embedded representation, we pass it through a projection layer to match the required dimensionality.
-We then concatenate it with the image and text embeddings and feed the combined representation into the transformer encoder.
-Appendix C Ensemble Methods
+A.1 Head-to-Fingertip Heatmap
+To generate this heatmap, we take the head and fingertip as reference points and construct a ray starting at the head and extending toward the image boundary in the fingertip direction.
+In the YouRefIt dataset [13], we use the detailed annotations provided in [13, 38], which include eye, elbow, wrist, and fingertip coordinates.
+These keypoints allow us to reliably define the reference direction for heatmap generation.
+In real-world settings where such annotations are unavailable, estimating these points is straightforward.
+For eye position, face detection followed by using the center of the detected face provides a strong approximation.
+Similarly, modern pose estimation models offer accurate wrist and fingertip predictions.
+For our experiments on the ISL Pointing and CAESAR datasets, we obtain necessary coordinates using OpenPose [9], which delivers robust and reliable keypoint detection.
+A.2 Wrist-to-Fingertip Heatmap
+We follow the same pipeline used for head-to-fingertip heatmap generation, but replace the eye coordinates with wrist coordinates and the fingertip coordinates remain the same.
+A.3 How to Create Heatmap
+We evaluate the Conic Attention Heatmap using and cone angles, as well as the Gaussian Ray Heatmap, which produces more localized activation.
+As shown in Tab.
+A.7, the Gaussian Ray Heatmap achieves the best performance.
+This suggests that broader attention regions may be less effective in practice, likely due to the high density of objects in typical scenes (see Fig.
+A.5).
+Therefore, larger heatmaps tend to cover multiple distractors, reducing the discriminative value of the spatial signal.
+| Heatmap style | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
+| Gaussian Ray Heatmap () | 72.9 | 62.2 | 35.1 | 0.2457 | 0.2490 |
+| Conic Attention Heatmap () | 70.0 | 60.0 | 33.9 | 0.2464 | 0.2632 |
+| Conic Attention Heatmap () | 71.5 | 59.4 | 30.8 | 0.2463 | 0.2764 |
+Appendix B Ensemble Methods
 In this section, we present the details of the ensemble methods we explored.
@@
 Similarly, we denote the confidence score of model as , and from model as .
-C.1 Confidence-Only
+B.1 Confidence-Only
 In the Confidence-Only approach, we focus solely on the top-1 prediction from each model.
@@
 These two predictions are then compared, and the one with the higher confidence score is selected as the final prediction.
-C.2 CLIP-Only (Top-1)
+B.2 CLIP-Only (Top-1)
 In this method, similar to the Confidence-Only approach, we select the bounding box with the highest confidence score from each model.
@@
 where is the embedded representation of cropped predicted bounding box of with the highest score.
-C.3 CLIP-Only (Top-2 + Threshold)
+B.3 CLIP-Only (Top-2 + Threshold)
 This method is similar to CLIP-Only (Top-1), where we take predictions from both models and compute the CLIP similarity score between the cropped bounding boxes and the input text.
@@
 Finally, we select the bounding box with the highest CLIP similarity score as the final prediction.
-The algorithm for this method is presented in Algorithm C.3.
-C.4 CLIP Fusion
+The algorithm for this method is presented in Algorithm B.3.
+B.4 CLIP Fusion
 In this method, we leverage both the CLIP similarity score and the model’s confidence score to make the final prediction.
@@
 Finally, we select the bounding box with the highest hybrid score as the final prediction.
-The algorithm for this method is presented in Algorithm C.4.
-C.5 CLIP-Aware Pointing Ensemble (CAPE)
+The algorithm for this method is presented in Algorithm B.4.
+B.5 CLIP-Aware Pointing Ensemble (CAPE)
 In this method, we combine the CLIP-Only (Top-2 + Threshold) and CLIP Fusion approaches.
@@
 For all other cases, we apply the CLIP-Only (Top-2 + Threshold) method.
-The algorithm is presented in Algorithm C.5.
+The algorithm is presented in Algorithm B.5.
+Appendix C Discussion
+C.1 Why is pointing heatmap useful?
+In ERU, the model must associate a textual instruction with the spatial region indicated by the person in the image.
+Without explicit spatial guidance, the model relies on implicit cues, such as hand position or object features, which become ambiguous when multiple candidate objects fit the text instruction.
+A pointing line provides a clear directional prior, narrowing the possible targets to regions intersecting or near the line.
+Formally, if is the set of candidate regions without explicit pointing heatmap (), then pointing guidance:
+where denotes the image, refers to the textual description, and is the pointing line. This reduces ambiguity and helps the model converge to the correct region.
+A second benefit is spatial alignment of multimodal embeddings.
+Without spatial cues, the latent representation of the intended object may blend with nearby distractors.
+The pointing line acts as a spatial attention prior, biasing the model toward regions consistent with the pointing direction.
+Finally, using a pointing heatmap also strengthens the gradient signal during training.
+The heatmap provides an additional, well-structured supervisory cue, reducing overfitting to spurious text or background correlations.
+Effectively, serves as a spatial inductive bias that stabilizes and accelerates learning.
+In summary, the pointing line provides an explicit spatial prior, constrains the candidate space, improves multimodal alignment, and yields more efficient and accurate grounding.
+C.2 Why does CAPE enhance the performance?
+The head-to-finger line provides a global directional prior, while the wrist-to-finger line captures local articulation.
+Together, they offer complementary spatial cues, allowing each network to capture different aspects of the pointing geometry and reducing the risk of misalignment across varied pointing poses.
+Therefore, we introduce CAPE module to ensemble the outputs of both models ( and ).
+We incorporate CLIP in addition to model confidence scores because CLIP learns a joint image–text embedding space where images and text with matching semantics are close.
+Formally, for an image region and a text description , CLIP provides a similarity score:
+where and are image and text encoders, and is cosine similarity. This score captures semantic alignment.
+Relying only on CLIP is insufficient because, while it provides semantic similarity, it does not encode geometric cues such as pointing direction.
+This can lead to imprecise localization in cluttered scenes or when multiple semantically plausible objects exist.
+Conversely, using only model confidence is also unreliable: pointing-based networks provide spatial priors but can be overconfident even when predictions are wrong due to occlusion, unusual poses, or network biases.
+The two pointing models provide spatial priors, while CLIP provides semantic priors.
+Combining them reduces both geometric and linguistic ambiguity.
+Conceptually, the fusion functions like a Bayesian ensemble.
+The pointing models supply a geometry-driven likelihood over regions, while CLIP acts as a semantic prior.
+Integrating these signals produces more sharper and reliable posterior estimate.
+As an alternative to CLIP model, we also experimented with SigLIP [80], which provides strong semantic grounding.
+However, replacing CLIP with SigLIP in the CAPE module yielded slightly lower performance (see Tab.
+C.8), so we continue using CLIP.
+In summary, CAPE fuses pointing-based confidence with CLIP similarity, integrating complementary spatial and semantic signals to improve referent prediction in challenging poses.
+| Method | IoU=0.25 | IoU=0.50 | IoU=0.75 |
+| CAPE with SigLIP | 74.5 | 65.1 | 35.4 |
+| CAPE with CLIP | 75.0 | 65.4 | 35.7 |
+| Method | Heatmap Input | IoU=0.25 | IoU=0.5 | IoU=0.75 |
+| Head-to-finger | 72.9 | 62.3 | 35.1 | |
+| Wrist-to-finger | 73.0 | 62.0 | 35.6 | |
+| Ground-Truth | 73.0 | 62.1 | 36.2 | |
+| None | 71.6 | 60.7 | 32.9 | |
+| Wrist-to-finger | 69.6 | 60.7 | 31.5 | |
+| Head-to-finger | 68.1 | 59.2 | 28.4 | |
+| Ground-Truth | 69.7 | 59.9 | 31.9 | |
+| None | 68.8 | 59.7 | 31.2 |
+C.3 Effect of Heatmap
+We evaluate the robustness of our models ( and ) using different heatmap inputs (Table C.9).
+First, we test using a wrist-to-finger heatmap.
+While IoU 0.25 and 0.75 show slight gains, integrating this into CAPE reduces overall performance, indicating a loss of complementary effectiveness.
+In our main experiments, we generate both heatmaps using predicted keypoints from a pose estimation model.
+In contrast, the Ground-Truth setup uses the annotated pointing coordinates, including the center of the ground-truth bounding box, to generate the heatmap.
+As expected, this yields a small performance boost since the line is guaranteed to intersect the target object’s center.
+However, the improvement is marginal, indicating that our predicted heatmaps are already accurate and that the pointing line reliably guides the model.
+Finally, providing a zero heatmap (None) removes explicit pointing guidance.
+Performance drops but remains reasonable, demonstrating that the model can rely on text and visual context.
+Notably, since the model is trained with explicit pointing heatmaps, it has learned to interpret pointing cues; therefore, even without a heatmap at inference, it can still achieve meaningful performance.
+Repeating the same experiments for shows similar trends, reinforcing the complementary nature of the two heatmaps.
+| Method | Text | IoU=0.25 | IoU=0.5 | IoU=0.75 |
+| Original | 72.9 | 62.3 | 35.1 | |
+| Dummy | 34.0 | 26.5 | 12.3 | |
+| Random | 24.9 | 19.0 | 9.4 | |
+| No Text | 39.2 | 27.9 | 10.2 | |
+| Original | 69.6 | 60.7 | 31.5 | |
+| Dummy | 34.6 | 25.5 | 11.0 | |
+| Random | 25.4 | 19.3 | 9.24 | |
+| No Text | 43.3 | 31.1 | 11.7 |
+C.4 Effect of Text Input
+In summary, for the embodied reference understanding task, text serves as the primary cue for object identification, while pointing information, derived from both the input image and the heatmap, plays a complementary, assistive role.
+In Table C.10, we evaluate our models under varying text input conditions to assess the role of textual guidance.
+The Original setup uses the dataset annotations.
+In the Dummy scenario, the text is replaced with the generic word “object,” while in the Random scenario, text is randomly sampled from other test examples, sometimes contradicting the actual target.
+Both scenarios cause a notable performance drop, highlighting the importance of textual input.
+Without accurate text instructions, the pointing line often intersects multiple objects, making it insufficient to identify the correct referent in cluttered scenes.
+In summary, text provides the primary cue for referent identification in ERU, while pointing information from the image and heatmap serves as a complementary signal.
+C.5 Finetuning or Frozen Text Encoder?
+We investigate the impact of freezing the text encoder during training.
+Based on multiple training experiments, we observed that finetuning the text encoder sometimes leads to instability, represented by the pink line in Fig.
+C.6.
+In contrast, freezing the text encoder results in consistently stable training, as shown by the green and orange lines in Fig.
+C.6.
+This indicates that finetuning introduces optimization sensitivity and may require multiple attempts to obtain a stable run.
+Despite this instability, finetuning ultimately yields better results.
+Table C.11 shows that, for both Setup A and Setup F, finetuning improves mAP at IoU 0.25 and 0.5, and also provides better scores.
+These gains suggest that adapting the text representations to the specific semantics of embodied reference understanding provides a meaningful advantage, even though the training process becomes less stable.
+| Text Encoder | IoU=0.25 | IoU=0.5 | IoU=0.75 | CLIP | |
+| Frozen, Setup A | 69.8 | 59.5 | 33.1 | 0.2454 | 0.2994 |
+| Finetuned, Setup A | 71.2 | 60.1 | 32.8 | 0.2469 | 0.2662 |
+| Frozen, Setup F | 71.3 | 62.1 | 33.7 | 0.2460 | 0.2634 |
+| Finetuned, Setup F | 72.9 | 62.2 | 35.1 | 0.2457 | 0.2490 |
 | Name | Value |
@@
 | Learning rate - text backbone | |
-| Learning rate | 1e-5 |
+| Learning rate | |
 | 2 | |
@@
 | Number of queries | 20 |
-Appendix D Finetuning or Frozen Text Encoder?
-Based on multiple training experiments, we observed that finetuning the text encoder sometimes leads to instability, represented by the pink line in Fig.
-E.5.
-In contrast, freezing the text encoder results in consistently stable training, as shown by the green and orange lines in Fig.
-E.5.
-Therefore, achieving stable training with a finetuned text encoder requires several attempts.
-However, despite this challenge, finetuning the text encoder ultimately yields better performance.
-Appendix E Hyperparameters and Other Training Parameters
-In Tab. C.8, we present empirically determined hyperparameters and other key training details to support reproducibility.
-Appendix F More Results
+Appendix D Hyperparameters and Other Training Parameters
 In Tab.
-F.9, we report the mAP at three different thresholds for all ablation setups, categorized by object size.
+C.12, we present empirically determined hyperparameters and other key training details to support reproducibility.
+For most hyperparameters, we follow established values from the literature, as they have been thoroughly analyzed in prior work [38, 11, 37, 43, 81, 85, 39].
+We also observe that training is not sensitive to the choice of loss coefficients, as the resulting accuracy differences are negligible.
+Appendix E More Results
+| IoU Threshold for mAP | 0.25 | 0.50 | 0.75 | |||||||||
+| Object Sizes | All | S | M | L | All | S | M | L | All | S | M | L |
+| PaliGemma2 | 54.4 | 46.2 | 53.0 | 64.1 | 51.2 | 42.4 | 47.2 | 64.1 | 44.5 | 26.2 | 43.3 | 64.1 |
+| Qwen2.5vl | 36.4 | 21.0 | 37.2 | 51.1 | 31.6 | 14.9 | 32.0 | 47.9 | 21.0 | 10.0 | 19.3 | 33.7 |
+| Grounding DINO | 47.6 | 40.7 | 47.9 | 54.4 | 46.8 | 38.7 | 47.4 | 54.4 | 44.5 | 33.1 | 46.1 | 54.4 |
+| Touch-Line-EWL | 50.7 | 49.8 | 53.2 | 45.6 | 46.5 | 45.5 | 49.3 | 41.3 | 24.6 | 20.3 | 30.9 | 29.3 |
+| Touch-Line-VTL | 45.9 | 44.2 | 51.2 | 32.6 | 40.3 | 39.5 | 43.2 | 32.6 | 20.6 | 13.9 | 29.4 | 32.6 |
+| 55.5 | 52.3 | 57.6 | 70.6 | 46.4 | 39.7 | 55.6 | 70.7 | 15.2 | 3.3 | 26.8 | 62.0 | |
+| 48.6 | 46.2 | 49.4 | 64.1 | 42.5 | 40.1 | 43.1 | 59.8 | 17.7 | 7.8 | 28.8 | 43.5 | |
+| CAPE w/ + | 55.5 | 52.9 | 55.8 | 75.0 | 50.5 | 47.3 | 50.9 | 75.0 | 18.6 | 9.1 | 26.6 | 57.6 |
+In Tab.
+E.13, we report the results of our model on the CAESAR dataset, which is simulation-based.
+We compare our model with SOTA methods, including Touch-Line, GroundingDINO, PaliGemma2, and Qwen2.5vl.
+Despite the significant domain gap, our model achieves very accurate performance, attaining SOTA scores in most cases.
+Under the IoU threshold of 0.75, PaliGemma2 and GroundingDINO achieve higher scores.
+In Tab.
+E.14, we report the mAP at three different thresholds for all ablation setups, categorized by object size.
 Similarly, in Tab.
-F.10, we present the CLIP scores and metrics for the same setups, also with respect to object size.
+E.15, we present the CLIP scores and metrics for the same setups, also with respect to object size.
 In Fig.
-F.6, we present examples of failure cases.
-In the first row, although our model correctly detects the target object, CAPE selects the prediction from , which is incorrect in this case.
-In the second row, the bottles are extremely difficult to distinguish, even with explicit pointing cues.
-Nevertheless, the predicted bounding boxes cover other bottles adjacent to the GT one, which is acceptable, as they overlap with the pointing line and are semantically consistent with the text.
-In the third row, both of our models, as well as the Touch-Line models, produce incorrect predictions.
-However, guided by the pointing information, the predictions from and lie on the pointing line and are very close to the GT object.
-In contrast, the Touch-Line predictions are entirely off the pointing line and significantly distant from the GT object.
+E.7, we present example failure cases from the ISL Pointing and CAESAR datasets.
+In the first column, all models predict the same object, as it aligns better with the pointing line and the text provides limited semantic guidance.
+Localization-related text is particularly challenging for the models to interpret.
+In the second column, only Touch-Line-VTL predicts the correct object.
+In the third column (from ISL Pointing), our models detect the scissors correctly, whereas Touch-Line-VTL detects only a small part of the spoon.
+Due to perspective and the lack of depth information, our models align the scissors with the pointing line.
+In Fig.
+E.8, we present additional example images from the YouRefIt dataset and qualitatively compare our model with SOTA models.
+CAPE achieves high accuracy in most cases, whereas Touch-Line often fails to detect the correct object, and the LMMs largely ignore the pointing cue, relying only on the text instruction.
+In Fig.
+E.9, we present example images from the ISL Pointing and CAESAR datasets.
+The first two columns show ISL Pointing images, while the remaining two columns contain CAESAR samples, which are simulation-based.
+In the first ISL example, all models except predict the object correctly.
+makes wrong prediction because its prediction aligns better with the wrist-to-fingertip pointing line than the GT object, and the provided text does not conflict with this prediction.
+In the second example, fails, but CAPE correctly selects ’s prediction.
+Touch-Line-VTL, PaliGemma2, Qwen2.5vl, and GroundingDINO make mistakes as they overlook the pointing cue.
+On the CAESAR dataset, despite the severe domain gap, our model and Touch-Line-VTL achieve strong performance.
+Our models closely follow the pointing cue and detect the correct object more accurately than other models.
+Given the challenging text instructions in both datasets, accurately and efficiently leveraging the pointing direction is key, explaining why CAPE surpasses all other models.
 | IoU Threshold for mAP | 0.25 | 0.50 | 0.75 | ||||||||||||
 | Setup | Object Sizes | Heatmap Injection | Ensemble | All | S | M | L | All | S | M | L | All | S | M | L |
+| PaliGemma2 | - | - | 58.8 | 29.0 | 53.5 | 75.8 | 46.9 | 22.1 | 50.8 | 68.0 | 31.7 | 6.2 | 34.1 | 54.8 | |
+| Qwen2.5vl-3b | - | - | 38.9 | 17.0 | 41.8 | 58.0 | 31.0 | 11.1 | 33.6 | 48.1 | 20.0 | 5.7 | 19.8 | 34.5 | |
+| Grounding DINO | - | - | 57.9 | 38.0 | 60.9 | 74.9 | 54.9 | 35.7 | 59.3 | 69.6 | 42.3 | 22.7 | 45.9 | 58.4 | |
 | Touch-Line-EWL [38] | - | - | 69.5 | 56.6 | 71.7 | 80.0 | 60.7 | 44.4 | 66.2 | 71.2 | 35.5 | 11.8 | 38.9 | 55.0 | |
@@
 | Setup | Object Sizes | Heatmap Injection | Ensemble | All | S | M | L | All | S | M | L |
-| Touch-Line-EWL [35] | - | - | 0.7369 | 0.6936 | 0.7305 | 0.7845 | 0.3168 | 0.3006 | 0.2903 | 0.3564 | |
-| Touch-Line-VTL [35] | - | - | 0.7368 | 0.6925 | 0.7321 | 0.7846 | 0.2843 | 0.2809 | 0.2276 | 0.3393 | |
-| 1 | Baseline | - | - | 0.7409 | 0.7003 | 0.7341 | 0.7861 | 0.2662 | 0.2182 | 0.2506 | 0.3266 |
-| 2 | Setup 1 + object center prediction | - | - | 0.7339 | 0.6947 | 0.7277 | 0.7787 | 0.2707 | 0.2201 | 0.2692 | 0.3222 |
-| 3 | Setup 1 + W2F heatmap | Embedding feature | - | 0.7353 | 0.6934 | 0.7297 | 0.7812 | 0.2984 | 0.2763 | 0.2616 | 0.3533 |
-| 4 | Setup 1 + H2F heatmap | Embedding feature | - | 0.7375 | 0.6936 | 0.7333 | 0.7852 | 0.2694 | 0.2649 | 0.2233 | 0.3172 |
-| 5 | Setup 2 + W2F heatmap | Embedding feature | - | 0.7344 | 0.6939 | 0.7306 | 0.7780 | 0.2770 | 0.2568 | 0.2319 | 0.3382 |
-| 6 | Setup 2 + H2F heatmap | Embedding feature | - | 0.7373 | 0.6953 | 0.7343 | 0.7819 | 0.249 | 0.222 | 0.2202 | 0.3024 |
-| 7 | Setup 1 + W2F heatmap + H2F heatmap | Embedding feature | - | 0.7345 | 0.6965 | 0.7277 | 0.7780 | 0.2744 | 0.2540 | 0.2314 | 0.3335 |
-| 8 | Setup 6 | Channel-wise input | - | 0.7348 | 0.6933 | 0.7273 | 0.7828 | 0.2959 | 0.2437 | 0.2642 | 0.3767 |
-| 9 | Setup 6 | Channel-wise feature | - | 0.7372 | 0.6953 | 0.7311 | 0.7839 | 0.2598 | 0.2289 | 0.2119 | 0.3339 |
-| 10 | Setup 5 + Setup 6 | Embedding feature | Confidence-Only | 0.7735 | 0.7441 | 0.7843 | 0.7923 | 0.247 | 0.2176 | 0.2042 | 0.3141 |
-| 11 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-1) | 0.7932 | 0.7688 | 0.8010 | 0.8098 | 0.2391 | 0.2242 | 0.2011 | 0.2878 |
-| 12 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-2 + Threshold) | 0.7379 | 0.7967 | 0.7343 | 0.7819 | 0.2422 | 0.2158 | 0.2009 | 0.3060 |
-| 13 | Setup 5 + Setup 6 | Embedding feature | CLIP Fusion | 0.7742 | 0.7927 | 0.7455 | 0.7842 | 0.2434 | 0.2137 | 0.2028 | 0.3090 |
-| 14 | Setup 5 + Setup 6 (Final Model) | Embedding feature | CAPE | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
-| 15 | Setup 6 w/ frozen text encoder | Embedding feature | - | 0.7382 | 0.6954 | 0.7323 | 0.7855 | 0.2634 | 0.2626 | 0.2233 | 0.3006 |
-| 16 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.7387 | 0.6962 | 0.7325 | 0.7860 | 0.2695 | 0.2685 | 0.2344 | 0.3028 |
-| 17 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.7391 | 0.6970 | 0.7332 | 0.7863 | 0.2765 | 0.2750 | 0.2499 | 0.3026 |
-| Ours - Final | Embedding feature | CAPE | 0.7983 | 0.7927 | 0.8015 | 0.8010 | 0.2476 | 0.2137 | 0.2241 | 0.3023 | |
+| Touch-Line-EWL [38] | - | - | 0.2456 | 0.2312 | 0.2435 | 0.2615 | 0.3168 | 0.3006 | 0.2903 | 0.3564 | |
+| Touch-Line-VTL [38] | - | - | 0.2456 | 0.2308 | 0.2440 | 0.2615 | 0.2843 | 0.2809 | 0.2276 | 0.3393 | |
+| 1 | Baseline | - | - | 0.2470 | 0.2334 | 0.2447 | 0.2620 | 0.2662 | 0.2182 | 0.2506 | 0.3266 |
+| 2 | Setup 1 + object center prediction | - | - | 0.2446 | 0.2316 | 0.2426 | 0.2596 | 0.2707 | 0.2201 | 0.2692 | 0.3222 |
+| 3 | Setup 1 + W2F heatmap | Embedding feature | - | 0.2451 | 0.2311 | 0.2432 | 0.2604 | 0.2984 | 0.2763 | 0.2616 | 0.3533 |
+| 4 | Setup 1 + H2F heatmap | Embedding feature | - | 0.2458 | 0.2312 | 0.2444 | 0.2617 | 0.2694 | 0.2649 | 0.2233 | 0.3172 |
+| 5 | Setup 2 + W2F heatmap | Embedding feature | - | 0.2448 | 0.2313 | 0.2435 | 0.2593 | 0.2770 | 0.2568 | 0.2319 | 0.3382 |
+| 6 | Setup 2 + H2F heatmap | Embedding feature | - | 0.2458 | 0.2318 | 0.2448 | 0.2606 | 0.249 | 0.222 | 0.2202 | 0.3024 |
+| 7 | Setup 1 + W2F heatmap + H2F heatmap | Embedding feature | - | 0.2448 | 0.2322 | 0.2426 | 0.2593 | 0.2744 | 0.2540 | 0.2314 | 0.3335 |
+| 8 | Setup 6 | Channel-wise input | - | 0.2449 | 0.2311 | 0.2424 | 0.2609 | 0.2959 | 0.2437 | 0.2642 | 0.3767 |
+| 9 | Setup 6 | Channel-wise feature | - | 0.2457 | 0.2318 | 0.2437 | 0.2613 | 0.2598 | 0.2289 | 0.2119 | 0.3339 |
+| 10 | Setup 5 + Setup 6 | Embedding feature | Confidence-Only | 0.2578 | 0.2480 | 0.2614 | 0.2641 | 0.247 | 0.2176 | 0.2042 | 0.3141 |
+| 11 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-1) | 0.2644 | 0.2563 | 0.2670 | 0.2699 | 0.2391 | 0.2242 | 0.2011 | 0.2878 |
+| 12 | Setup 5 + Setup 6 | Embedding feature | CLIP-Only (Top-2 + Threshold) | 0.2460 | 0.2656 | 0.2448 | 0.2606 | 0.2422 | 0.2158 | 0.2009 | 0.3060 |
+| 13 | Setup 5 + Setup 6 | Embedding feature | CLIP Fusion | 0.2581 | 0.2642 | 0.2485 | 0.2614 | 0.2434 | 0.2137 | 0.2028 | 0.3090 |
+| 14 | Setup 5 + Setup 6 (Final Model) | Embedding feature | CAPE | 0.2661 | 0.2642 | 0.2672 | 0.2670 | 0.2476 | 0.2137 | 0.2241 | 0.3023 |
+| 15 | Setup 6 w/ frozen text encoder | Embedding feature | - | 0.2461 | 0.2318 | 0.2441 | 0.2618 | 0.2634 | 0.2626 | 0.2233 | 0.3006 |
+| 16 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.2462 | 0.2321 | 0.2442 | 0.2620 | 0.2695 | 0.2685 | 0.2344 | 0.3028 |
+| 17 | Setup 6 w/ Conic Attention Heatmap () | Embedding feature | - | 0.2464 | 0.2323 | 0.2444 | 0.2621 | 0.2765 | 0.2750 | 0.2499 | 0.3026 |
+| Ours - Final | Embedding feature | CAPE | 0.2661 | 0.2642 | 0.2672 | 0.2670 | 0.2476 | 0.2137 | 0.2241 | 0.3023 | |